### v1

# ProLLaMA: A Protein Large Language Model for Multi-Task Protein Language Processing

Abstract
Large Language Models (LLMs), including GPT-x and LLaMA2, have achieved remarkable performance in multiple Natural Language Processing (NLP) tasks. Under the premise that protein sequences constitute the protein language, Protein Large Language Models (ProLLMs) trained on protein corpora excel at de novo protein sequence generation. However, as of now, unlike LLMs in NLP, no ProLLM is capable of multiple tasks in the Protein Language Processing (PLP) field. This prompts us to delineate the inherent limitations in current ProLLMs: (i) the lack of natural language capabilities, (ii) insufficient instruction understanding, and (iii) high training resource demands. To address these challenges, we introduce a training framework to transform any general LLM into a ProLLM capable of handling multiple PLP tasks. Specifically, our framework utilizes low-rank adaptation and employs a two-stage training approach, and it is distinguished by its universality, low overhead, and scalability. Through training under this framework, we propose the ProLLaMA model, the first known ProLLM to handle multiple PLP tasks simultaneously. Experiments show that ProLLaMA achieves state-of-the-art results in the unconditional protein sequence generation task. In the controllable protein sequence generation task, ProLLaMA can design novel proteins with desired functionalities. In the protein property prediction task, ProLLaMA achieves nearly 100% accuracy across many categories. The latter two tasks are beyond the reach of other ProLLMs. Code is available at https://github.com/Lyu6PosHao/ProLLaMA.
Machine Learning, ICML
1 Introduction
Large Language Models (LLMs), like GPT-x and LLaMA2 (Bubeck et al., 2023; Touvron et al., 2023), have achieved outstanding performance in handling a wide range of Natural Language Processing (NLP) tasks (Tamkin et al., 2021; Zhao et al., 2023; Kocoń et al., 2023; Qin et al., 2023; Huang et al., 2023; Zhong et al., 2023; Bang et al., 2023), including both Natural Language Generation (NLG) and Natural Language Understanding (NLU) tasks. This surge in LLMs has extended their applications beyond traditional contexts, including their adoption in the challenging field of protein design (Notin et al., 2022; Strokach & Kim, 2022; Ferruz & Höcker, 2022).
Figure 1: Analogous to general LLMs managing multiple tasks in the NLP field, our ProLLaMA can handle multiple tasks in Protein Language Processing. In contrast, other ProLLMs are limited to the single task of unconditional protein generation.
Taking protein sequences as the protein language, researchers train models with architectures similar to current LLMs on a vast protein corpus (Strokach & Kim, 2022; Ferruz et al., 2022; Nijkamp et al., 2023; Moffat et al., 2022; Madani et al., 2023; Qin et al., 2023). These Protein LLMs (ProLLMs) have paved the way for the rapid generation of structurally plausible protein sequences, holding immense potential for biomedical and biotechnological innovations (Notin et al., 2022; Nijkamp et al., 2023). However, this progress is met with challenges, particularly in extending their capabilities beyond sequence generation.
Analogous to NLP, tasks related to protein language can be viewed as Protein Language Processing (PLP) (Bepler & Berger, 2021; Ofer et al., 2021). It is evident that current ProLLMs primarily focus on a single task in PLP, not covering multiple tasks like LLMs do in NLP (Ferruz et al., 2022; Madani et al., 2023). These limitations prompt the need for innovative solutions to unleash the full potential of ProLLMs. Developing a multi-tasking ProLLM would be highly beneficial for protein engineering and understanding the protein fitness landscape (Wright et al., 1932; Pan & Kortemme, 2021; Ren et al., 2022; Song & Li, 2023), but three main challenges must be considered:
(i) Necessity of Natural Language: Protein language is not fully sufficient for PLP tasks, meaning it cannot fully represent all components of a task (the user instruction and the expected output) (Xu et al., 2023; Wang et al., 2023). The instruction and the output require a language beyond protein language (typically, natural language) for representation, which current ProLLMs lack.
(ii) Instruction Following: To possess multi-tasking capabilities, models must execute tasks following user instructions (Zeng et al., 2023; Liu et al., 2023b; Zhou et al., 2023). However, current ProLLMs are unable to follow instructions.
(iii) Training Resource Consumption: Substantial training resources are needed for models to learn natural language, protein language, and user instructions (Cui et al., 2023), which can sometimes be unaffordable.
To address the challenges, we propose a two-stage training framework to achieve a ProLLM for multi-task PLP. In the first stage, we leverage a pre-trained general LLM like LLaMA2 to continually learn the protein language while maintaining the natural language knowledge of the model. In the second stage, the model is further trained on a multi-task PLP dataset through instruction tuning to equip the model with instruction following capabilities and multi-task capabilities of PLP. Notably, during both stages, we adopt Low-Rank Adaptation (LoRA) (Hu et al., 2021), which prevents catastrophic forgetting in the first stage, enhances scalability in the second stage and reduces the training cost.
Using the training framework, we develop ProLLaMA, a model capable of multi-tasks for PLP, distinguishing it from all other ProLLMs. Through a series of experiments, we demonstrate the multitasking capabilities of our ProLLaMA. Specifically, in unconditional protein generation, ProLLaMA outperforms current ProLLMs and other sequence-generative models on common metrics such as pLDDT and TM-score. In controllable protein generation, based on a user-provided textual description of the desired protein functionalities, ProLLaMA generates novel proteins from scratch with functionalities as competent as natural proteins, such as SAM-MT and Trx. For protein property prediction, ProLLaMA achieves an average accuracy of 72% on the test dataset and nearly 100% accuracy in many subcategories. In summary, the contributions of our research are as follows:
$\bullet$ We propose a training framework with scalability and efficiency that enables any general LLM to be trained as a proficient model for multiple tasks in Protein Language Processing.
$\bullet$ Our ProLLaMA is the first model to our knowledge capable of simultaneously handling multiple PLP tasks.
$\bullet$ Experiments show that our ProLLaMA not only handles PLP tasks beyond the reach of existing ProLLMs but also achieves state-of-the-art results in the protein generation task where current ProLLMs are active.
2 Preliminaries
2.1 Necessity of Natural Language
As aforementioned, given the similarities between protein sequences and natural language, tasks related to protein sequences can be considered as Protein Language Processing (PLP), analogous to NLP. However, we observe a fundamental difference between protein language and natural language: natural language is complete for NLP tasks, whereas protein language is not complete for PLP tasks.
Figure 2: The overall architecture of our ProLLaMA. We add low-rank adapters (LoRA) to certain weights. During the training process, we freeze these weights and other parameters, focusing solely on training LoRA.
Specifically, natural language can represent all components of a task (i.e., the input instructions and the expected output) (Liu et al., 2023a). For example, in a sentiment analysis task, the instruction and output can be expressed as “Analyze the sentiment contained in this sentence: I am happy” and “The sentiment is positive” respectively, all in natural language. However, for PLP tasks, the instructions and outputs cannot be fully represented in protein language. For instance, in the protein property prediction task, the task instruction could be ”Predict the property of this protein: MAFCF…FEV”, with the expected output being ”The property is Trx superfamily.” It is evident that both the instructions and outputs of tasks require assistance from a language beyond protein language, in this case, natural language, for representation.
Therefore, multi-task ProLLMs must possess a certain level of natural language ability, especially as more textual descriptions of proteins become available (Xu et al., 2023).
2.2 Causal Language Modeling
Causal Language Modeling (CLM) is the objective of training LLMs. Given a token sequence $\boldsymbol{x}=\{x_{0},x_{1},\ldots,x_{n-1}\}$ that is subject to training. CLM can be conceptualized as predicting the $i$-th token based on the preceding $i-1$ tokens (Bengio et al., 2000). Therefore, the optimization objective when training LLM using CLM is formulated as:
$\small\mathcal{L}(\Theta)=\mathbb{E}_{\boldsymbol{x}\sim\mathcal{D}}\left[-\sum_{i}\log p(x_{i}|x_{0},x_{1},\ldots,x_{i-1};\Theta)\right]$ (1)
where $\mathcal{L}$ denotes the loss function, $\Theta$ denotes the trainable parameters of the model, and $\mathcal{D}$ is the training dataset.
2.3 Low Rank Adaptation
Low-Rank Adaptation (LoRA) (Hu et al., 2021) is a parameter-efficient technique for fine-tuning LLMs. Due to the immense parameter size of LLMs, full-parameter fine-tuning could be impractical sometimes. LoRA circumvents this by freezing the original parameters of LLMs and introducing additional trainable low-rank adapters. It achieves fine-tuning with a significantly smaller number of trainable parameters, yielding results comparable to full-parameter fine-tuning.
[…]
Consequently, the quantity of parameters involved in training is reduced from $dh$ to $r(d+h)$. Given the low-rank hypothesis, where $r\ll d$ and $r\ll h$, this reduction in the number of trainable parameters is quite significant.
Figure 3: The overall training framework of our ProLLaMA. (A-B): Training is divided into two stages. (A): By continual learning on UniRef50, LLaMA2 learns protein language, resulting in ProLLaMA. (B) By instruction tuning on the instruction dataset, ProLLaMA is capable of multi-tasking. (C) By instruction tuning again on your own dataset, ProLLaMA can be expanded to more tasks.
During training, the original parameters of the model are frozen, with only the low-rank adapters trainable. After training, the newly acquired knowledge is stored in the low-rank adapters, namely A and B. Additionally, A and B can be integrated into the original model using Equation 2. Although this integration means the low-rank adapters are no longer plug-and-play, it ensures that the architecture of the post-training model remains consistent with that of the original model. Finally, LoRA prevents catastrophic forgetting of the original knowledge, as the newly learned knowledge has a lower rank than the original knowledge.
3 Methods
The overview of the architecture of our ProLLaMA is shown in Figure 2, and the overview of our training framework is shown in Figure 3. In Section 3.1, we show how the LLaMA2 model learns protein language through continual learning, resulting in ProLLaMA. In Section 3.2, we show the integration of various tasks into ProLLaMA through instruction tuning. In Section 3.3, we show more tasks can be integrated into ProLLaMA. These three sections correspond to Figure 3(A-C).
3.1 Continual Learning on Protein Language
As mentioned in Section 1, current ProLLMs are lack of natural language abilities, which hinders multi-task capabilities. To solve this problem, we propose leveraging a pre-trained LLaMA2 to perform continual learning on protein language, as shown in Figure 3(A). This approach is analogous to humans learning a foreign language, where the model learns protein language while retaining its original natural language abilities.
Specifically, we construct a dataset based on UniRef50 (Suzek et al., 2015). We preprocess each protein sequence with specific prefixes and suffixes. This standardized format aids LLaMA2 in distinguishing the new protein language from its existing natural language knowledge, thereby reducing confusion.
[…]
We train the model with Equation 1 as the target on the aforementioned dataset, resulting in ProLLaMA. Benefiting from LoRA, we train only 10% of the parameters, in contrast to full-parameter training, which significantly reduces training costs. Additionally, as the remaining parameters are not involved in training, the inherent natural language abilities of the model are preserved.
In summary, by enabling a pre-trained LLaMA2 to continually learn protein language and utilizing LoRA, we have developed ProLLaMA, a model that comprehends both protein sequence and natural language. Consequently, we have addressed the problem mentioned in the introduction: the lack of natural language abilities and excessive consumption of training resources.
3.2 Performing Various Tasks
As mentioned in Section 1, current ProLLMs are unable to perform multiple tasks based on user instructions. To solve this problem, we perform instruction tuning on the ProLLaMA obtained from the previous section, as shown in Figure 3(B).
To be specific, we construct a multi-task dataset, where each item represents a task sample. A task consists of two parts: the instruction and the output. Training conducted on this dataset can be denoted as follows:
$\mathcal{L}(\Theta)=\mathbb{E}_{\boldsymbol{x,u}\sim\mathcal{D}}~{}[-\log~{}p(\boldsymbol{x}|\boldsymbol{u};\Theta)]$ (4)
Here, $\Theta$ denotes the parameters to be optimized, $\mathcal{L}$ the loss function, and $\mathcal{D}$ the dataset. $\boldsymbol{u}$ denotes the instruction of a task. $\boldsymbol{x}=\{x_{0},x_{1},\ldots,x_{n-1}\}$ denotes the output of a task, where $x_{i}$ is the $i$-th token of the output.
Since ProLLaMA employs causal language modeling, we need to combine Equation 4 with Equation 1:
$\small\mathcal{L}(\Theta)=\mathbb{E}_{\boldsymbol{x}\sim\mathcal{D}}\left[-\sum_{i}\log p(x_{i}|\boldsymbol{u},x_{0},x_{1},\ldots,x_{i-1};\Theta)\right]$ (5)
Equation 5 is the optimization objective for instruction tuning of ProLLaMA. $\boldsymbol{u}$ is not involved in the loss calculation, whereas $\boldsymbol{x}$ is. This is because the latter is the output part, where ensuring its quality of generation is crucial. The former, $\boldsymbol{u}$, as an instruction signal, only needs to be understood by the model and does not require generation. In the instruction tuning stage, we exclusively train LoRA at a lower rank than specified in Section 3.1.
Table 1: Comparison of proteins generated by different models. Our ProLLaMA achieves the best performance on pLDDT, TM-score, and RMSD metrics, and is second-best in SC-Perp, demonstrating ProLLaMA excels in de novo protein design.
Architecture Method pLDDTnormal-$\uparrow$ SC-Perpnormal-$\downarrow$ AFDB PDB
TM-scorenormal-$\uparrow$ RMSDnormal-$\downarrow$ TM-scorenormal-$\uparrow$ RMSDnormal-$\downarrow$
CNN CARP (Alamdari et al., 2023) 34.40±14.43 4.05±0.52 0.28 19.38 0.38 8.95
LRAR (Alamdari et al., 2023) 49.13±15.50 3.59±0.54 0.40 14.47 0.43 9.47
AutoEncoder ESM-1b (Rives et al., 2021) 59.57±15.36 3.47±0.68 0.34 20.88 0.44 8.59
ESM-2 (Lin et al., 2023) 51.16±15.52 3.58±0.69 0.20 35.70 0.41 9.57
Diffusion
EvoDiff (Alamdari et al., 2023) 44.29±14.51 3.71±0.52 0.32 21.02 0.41 10.11
LLM ProtGPT2 (Ferruz et al., 2022) 56.32±16.05 3.27±0.59 0.44 12.60 0.43 9.19
ProGen2 (Nijkamp et al., 2023) 61.07±18.45 2.90±0.71 0.43 15.52 0.44 11.02
ProLLaMA (ours) 66.49±12.61 3.10±0.65 0.49 9.50 0.48 7.63
In summary, through instruction tuning, we have made ProLLaMA capable of following instructions and performing multiple tasks. In contrast, other ProLLMs remain limited to the single task of protein sequence generation. Consequently, we have addressed the problems mentioned in the introduction: the lack of instruction following and the lack of multi-task capabilities.
3.3 Expanding to More Tasks
In this section, we demonstrate that benefiting from the scalability of the training framework, ProLLaMA can be easily extended to a broader range of tasks as shown in Figure 3(C).
To be specific, to adapt ProLLaMA to more tasks, researchers must customize the instruction dataset and then perform instruction tuning on ProLLaMA again with this dataset. ProLLaMA can be either the one obtained in the first stage or the one obtained in the second stage. Benefiting from LoRA, such instruction tuning requires only minimal training resources. An instance is shown in Figure 3(C). The new task in this instance is protein complex prediction, where the model must ascertain whether two input proteins can form a complex. In this task, the instruction $\boldsymbol{u}$ comprises a sentence describing the user intent and two protein sequences, while the output $\boldsymbol{x}$ should be a judgment of Yes or No.
Table 2: Controllable generation on four different instructions. Given SAM-MT, TPHD, Trx, and CheY as instructions, ProLLaMA generates proteins with the desired functionalities. High values of TM-score and H-Prob indicate the generated proteins meet the instructions, in contrast to the uncontrollable generation of other models.
Method SAM-MT TPHD Trx CheY
TM-scorenormal-$\uparrow$ H-Probnormal-$\uparrow$ TM-scorenormal-$\uparrow$ H-Probnormal-$\uparrow$ TM-scorenormal-$\uparrow$ H-Probnormal-$\uparrow$ TM-scorenormal-$\uparrow$ H-Probnormal-$\uparrow$
ESM-1b 0.58 0.37 0.55 0.48 0.61 0.37 0.63 0.27
ESM-2 0.52 0.26 0.51 0.25 0.53 0.30 0.57 0.18
EvoDiff 0.46 1.17 0.42 1.80 0.42 1.10 0.46 1.43
ProtGPT2 0.45 3.86 0.43 4.62 0.44 2.53 0.45 4.86
ProGen2 0.44 1.90 0.45 2.49 0.43 2.44 0.44 2.13
ProLLaMA (ours) 0.71 98.13 0.82 100.00 0.93 99.96 0.81 100.00
In summary, our framework and model are flexible and easily extensible. With collaborative efforts from other researchers, ProLLaMA has the potential to become more powerful.
4 Experiments
We introduce the experiment setup in Section 4.1. And we evaluate the unconditional protein generation task in Section 4.2, the controllable protein generation task in Section 4.3, the protein property prediction task in Section 4.4, and the natural language ability in Section 4.5.
4.1 Experiment Setup
Training Settings: For continual learning, the LoRA rank is set to 128, employing the AdamW optimizer alongside a cosine annealing scheduler with warm-up. The peak learning rate stands at 0.05, with a total of one training epoch. For instruction tuning, the LoRA rank is set to 64 with two training epochs, and all other settings remain consistent with the continual learning setup.
Training Datasets: For continual learning, the dataset used consists of protein sequences from UniRef50 (Suzek et al., 2015). For instruction tuning, the instruction dataset is constructed using protein sequences from UniRef50 and protein property texts from InterPro (Paysan-Lafosse et al., 2023).
Evaluation Settings: Unconditional protein generation involves generating protein sequences without specific instructions. Controllable protein generation involves generating desired protein sequences based on instructions that specify the required functionalities. Property prediction involves predicting protein properties based on instructions, which include the protein sequences to be predicted.
Evaluation Metrics: We use the following metrics to evaluate the generated protein sequences. The pLDDT (Jumper et al., 2021) is used to measure whether sequences are structurally plausible. Self-Consistency Perplexity (SC-Perp) (Alamdari et al., 2023) serves as an additional metric of plausible structures since pLDDT falls short in dealing with intrinsically disordered regions (IDRs) (Davey, 2019). TM-score (Zhang & Skolnick, 2004) reflects the structural similarity between generated sequences and known ones in AFDB (Varadi et al., 2022) and PDB (Berman et al., 2002). RMSD also reflects the structural similarity from the perspective of atomic distance. Homologous probability (H-Prob) reflects the probability that the generated protein is homologous to a known one. Seq-Ident reflects the sequence similarity between generated sequences and known ones.
4.2 Unconditional Protein Generation
We compare our model with other state-of-the-art models in the protein sequence generation field. These models encompass a wide range of architectures and are trained exclusively on protein sequences. Table 1 shows the results. Our ProLLaMA is optimal on pLDDT, TM-score, and RMSD and is suboptimal on SC-Perp. This indicates that ProLLaMA, through its training on protein sequence data, can generate structurally plausible proteins. Additionally, it is important to note that ProLLaMA-generated proteins exhibit a mean and standard deviation for pLDDT and SC-Perp of 66.49±12.61 and 3.10±0.65, respectively. These values are comparable to those of natural proteins as reported by Alamdari et al. (2023), which are 68.25±17.85 and 3.09±0.63, respectively.
Figure 4: Contrast experiment of ProLLaMA. (A-C): Compared to the baseline (ESM2), ProLLaMA maintains a high, and even higher, quality of generated proteins as their length increases. (D-F): For the four superfamilies of CheY, TPHD, Trx, and SAM-MT (SAM), the proteins generated by ProLLaMA are comparable to natural proteins.
De novo design of long and structurally plausible protein sequences is highly challenging (Ferruz et al., 2022), yet our ProLLaMA excels in meeting this challenge. Figure 4(A) shows that as the sequence length increases, the pLDDT of proteins generated by ProLLaMA does not decrease but rather increases. In contrast, the pLDDT of proteins generated by ESM2 significantly decreases as the length increases. This indicates that ProLLaMA is able to capture long-range dependencies between amino acids, while ESM2 faces difficulties in the de novo design of long protein sequences. Figure 4(B) and Figure 4(C) also support this point. Figure 4(B) shows that as the length increases, SC-Perp of ProLLaMA slightly increases and then continues to decrease, while SC-Perp of ESM2 does the opposite. Figure 4(C) shows that as the length increases, the TM-score of ESM2 gradually decreases from above 0.6 to less than 0.2. In contrast, the TM-score of ProLLaMA remains above 0.5 for nearly all lengths, even exceeding 0.8 when the length is greater than 350. These demonstrate the robust sequence generation capability of ProLLaMA, especially in generating longer sequences.
4.3 Controllable Protein Generation
To instruct ProLLaMA in controllable protein generation, we utilize four superfamily descriptions as instructions respectively: the S-adenosyl-L-methionine-dependent methyltransferase superfamily (SAM-MT), the Tetratricopeptide-like helical domain superfamily (TPHD), the Thioredoxin-like superfamily (Trx), and the CheY-like superfamily (CheY). For each superfamily, ProLLaMA generates 100 protein sequences. Additionally, we use 100 natural proteins from each of the four superfamilies as benchmarks for comparison. We employ Foldseek to compare the generated proteins with the natural ones.
Table 3: Protein property prediction. Our ProLLaMA achieves nearly 100% accuracy for many superfamilies. Here are 10 of them.
OBFD UPF0145 NACD U3S CCHC Kazal SAM-MT TPHD Trx CheY
Accuracy 100.00% 100.00% 100.00% 100.00% 95.24% 100.00% 93.67% 90.84% 94.17% 100.00%
The results shown in Table 2 demonstrate that ProLLaMA can generate desired protein sequences based on instructions that specify the required functionalities, confirming the capability for controllable generation. For SAM-MT, the TM-scores of our generated sequences exceed 0.7; for TPHD and CheY, they are over 0.8; and for Trx, they surpass 0.9. The high TM-score indicates that the structures of the generated proteins closely resemble those of natural proteins in the same superfamily, implying functional similarity. For SAM-MT, TPHD, Trx, and CheY, all of the H-prob values are close to or even equal to 100%, indicating that the generated proteins are homologous to natural proteins and belong to the same superfamily. In summary, these provide strong evidence that the protein generation of ProLLaMA is controllable under instructions. In contrast, other models exhibit low TM-score and very low H-Prob due to their uncontrollable generation.
Additionally, using natural proteins as a benchmark, we assess pLDDT, SC-Perp, and TM-score of proteins generated by ProLLaMA. Figure 4(D) shows that for CheY, TPHD, Trx, and SAM-MT, the average pLDDT of generated proteins is only 19.0%, 1.41%, 10.9%, and 10.3% lower than that of natural proteins, respectively. Figure 4(E) shows that for TPHD and SAM-MT, the average SC-Perp is 5.66% and 5.06% lower; for Trx and CheY, the average SC-Perp is 7.58% and 14.92% higher. Figure 4(F) visualizes the TM-score, with scores near the maximum indicating a high degree of structural similarity. These findings indicate that proteins generated by ProLLaMA are comparable to their natural counterparts in the same superfamily. Given that averages are considered, there are instances where generated proteins outperform natural proteins.
Figure 5: Protein visualization. Cases of controllable generation by ProLLaMA using SAM-MT, TPHD, Trx, and CheY as instructions. Blue is generated proteins, and yellow is natural. They are similar in structure and function but different in sequence.
In Figure 5, we present actual examples of proteins generated by ProLLaMA (depicted in blue) alongside the most structurally similar natural proteins from PDB (depicted in yellow). The four proteins are generated under the instructions of SAM-MT, TPHD, Trx, and CheY. We find that the four natural proteins also belong to SAM-MT, TPHD, Trx, and CheY, respectively, according to information from InterPro. This implies functional similarity between the generated proteins and natural ones, which further validates the effectiveness of controllable generation. The significant overlap in the 3D structures and the high TM-score confirm structural similarity. The lower Seq-Ident scores indicate sequence diversity. In summary, through controllable protein generation, ProLLaMA is capable of generating desired proteins with functions and structures similar to natural proteins, yet with novel sequences.
4.4 Property Prediction
We use the test dataset to evaluate whether ProLLaMA can predict the superfamily to which a given protein belongs. The test dataset consists of 10,000 samples. One protein may belong to two or more superfamilies simultaneously. Therefore, for the i-th protein in the test dataset, we denote its true superfamily text description as the set $F_{i}$ and the prediction by ProLLaMA as the set $F^{\prime}_{i}$. We calculate the prediction accuracy using:
$\text{Acc}=\dfrac{\sum\nolimits_{i=1}^{N}\left|F_{i}\cap F^{\prime}_{i}\right|}{\sum\nolimits_{i=1}^{N}\left|F^{\prime}_{i}\right|}$ (6)
where $\left|X\right|$ denotes the number of elements in set $X$, $\cap$ denotes the intersection of two sets, and $N$ denotes the number of samples in the test dataset.
Although ProLLaMA actually performs a classification task here, it is more complex than typical ones. The key difference is that typical classification tasks require models to output a fixed label, often in one-hot encoding. In contrast, ProLLaMA outputs a full textual description of the result, significantly increasing the challenge. Even so, ProLLaMA still achieves an average accuracy of 72% on the test dataset. In addition, ProLLaMA achieves considerably high accuracy in many specific superfamily categories. Table 3 lists ten of them, showing that the accuracy exceeds 90% in all the ten superfamilies and reaches 100% in OBFD, UPF0145, NACD, U3S, Kazal, and CheY.
4.5 Natural Language Ability
Table 4: Comparison of the natural language ability. Vocab represents whether the vocabulary of the model supports natural language. Generation represents the sentence generation.
Type Model Vocab Generation QA
GeneralLLM LLaMA2 45% 44%
ProLLM ProGen2 - -
ProtGPT2 0% -
ProLLaMA 26% 33%
We evaluate the natural language ability of ProLLaMA and other models. As shown in Table 4, the vocabulary of ProGen2 only includes uppercase letters representing amino acids, failing to cover all English words and phrases. As for the sentence generation task on Wikipedia text, the accuracy of ProtGPT2 is 0%, indicating that it lacks natural language abilities. The accuracy of ProLLaMA is 26%, compared to 45% of LLaMA2. As for the QA task, the accuracy of ProLLaMA is 33%, compared to 44% of LLaMA2. These suggest that ProLLaMA has natural language abilities, albeit at a reduced level compared to LLaMA2.
5 Related Work
Protein Language Models
ProLLMs aforementioned are a subset of Protein Language Models (PLMs). Recognizing the similarity between natural language sequences and protein sequences, many algorithms from NLP have been applied to protein sequence data (Yang et al., 2018; Alley et al., 2019; Rao et al., 2021; Elnaggar et al., 2021). This has led to the development of PLMs, which are broadly categorized into two types (Ferruz & Höcker, 2022; Zheng et al., 2023).
One type adopts the decoder-only architecture and Causal Language Modeling (CLM) (Bengio et al., 2000; Vaswani et al., 2017), similar to general LLMs. Therefore, we refer to it as the Protein Large Language Model (ProLLM). ProLLMs primarily concentrate on de novo protein sequence generation (Moffat et al., 2022; Ferruz et al., 2022; Madani et al., 2023; Nijkamp et al., 2023), with a minority also focusing on fitness prediction (Notin et al., 2022). The other type adopts the encoder-only architecture and Masked Language Modeling (MLM) (Devlin et al., 2018; Meier et al., 2021; Rives et al., 2021; Brandes et al., 2022; Lin et al., 2023). They excel in protein representation learning, with the learned representations being applied to downstream tasks such as property prediction (Xu et al., 2023). However, they face challenges in de novo protein design.
Our ProLLaMA is capable of multitasking, excelling in tasks that both of the above types specialize in, surpassing existing ProLLMs and even PLMs.
Training LLMs
There is a well-established procedure for training LLMs (Zhao et al., 2023). Initially, models undergo pre-training on large-scale, unlabeled corpora to grasp grammatical and semantic principles (Min et al., 2023). Then, models are subjected to instruction tuning using an instruction dataset, enabling models to understand instructions and perform various tasks (Zhang et al., 2023). However, current ProLLMs cannot undergo instruction tuning due to the lack of natural language abilities.
Various parameter-efficient techniques have been proposed to accelerate training and conserve memory (He et al., 2021; Li & Liang, 2021; Hu et al., 2021; Liu et al., 2022; Hu et al., 2023), like Low-rank Adaptation (LoRA). We propose the use of LoRA in our training framework, which prevents catastrophic forgetting, enhances scalability, and reduces training costs. To our knowledge, this is the first application of parameter-efficient training techniques for ProLLMs.
6 Conclusion
Existing ProLLMs excel in the single task of protein generation but fall short in multiple tasks. In this work, we introduce an efficient training framework to transform any general LLM into a multi-task ProLLM. We also develop ProLLaMA, a versatile ProLLM for multiple tasks like controllable protein generation and protein property prediction. Experiments indicate that ProLLaMA performs exceptionally well. We are confident that our framework and model will have a significant impact on the AI4Science community.
Impact Statements
Our work holds the potential to revolutionize the field of computational biology and biotechnology. With multi-task capabilities in the protein field, our ProLLaMA could significantly accelerate the pace of research and innovation in areas such as drug discovery, synthetic biology, and the development of novel biomaterials.
Moreover, our work can enable a wider range of researchers and institutions, particularly those with limited resources, to participate in cutting-edge research, fostering a more inclusive AI4Science community.
However, our work may raise concerns regarding model reliability and the potential for misuse. The uncertainty in model outputs could lead to misinformed decisions in critical research areas like drug discovery, necessitating cautious reliance and thorough validation by human experts to mitigate risks. Moreover, the powerful capabilities of ProLLaMA could be exploited for harmful purposes, posing bio-security challenges and ethical challenges.
A collaborative effort among researchers, policymakers, and the broader community is crucial to harness the benefits of our work while addressing the potential risks and ethical considerations.
Appendix A Appendix
A.1 Detailed Training Settings
For continual learning, the block size is set to 2048, with a gradient accumulation step count of 8. The warm-up ratio is 0.05, and the weight decay is set to 0.01. The data type is bfloat16. The batch size per GPU is 4. DeepSpeed Zero Redundancy Optimizer stage 2 (ZeRO-2) is employed without using offload. For instruction tuning, the max sequence length is set to 256, with a gradient accumulation step count of 4. The warm-up ratio is 0.03. The batch size per GPU is 144. Other settings are the same as continual learning.
A.2 Detailed Evaluation Metrics
We use OmegaFold to calculate the values of the predicted Local Distance Difference Test (pLDDT) of protein sequences. Omegafold performs structure prediction without the need for homologous sequences or evolutionary information, relying solely on a single sequence for prediction. To calculate Self-Consistency Perplexity (SC-Perp), following the process mentioned in EvoDiff, we fold the sequence into a 3D structure using OmegaFold, then unfold it back into the sequence using ProteinMPNN. The self-consistency perplexity between the resulting sequence and the original sequence is referred to as SC-Perp.
We calculate TM-score, RMSD, H-Prob, and Seq-Ident using Foldseek. Foldseek facilitates the pairing of the queried protein $p^{query}$ with structurally similar proteins from an existing protein database (AFDB or PDB), yielding pairs represented as $(p^{query},\ p^{target})$. Here, $p^{target}$ denotes the protein in the database with a significant structural similarity to $p^{query}$. The magnitude of the average Template Modeling score (TM-score) value and Root-Mean-Square Deviation (RMSD) reflects the degree of structural similarity. TM-score takes into account the overall topological structure of proteins, focusing more on the overall structure. RMSD calculates the square root of the average position deviation of corresponding atoms between two protein structures, being highly sensitive to the size of the protein structure and local variations. Additionally, Foldseek also calculates the Sequence Identity (Seq-Ident) between $p^{query}$ and $p^{target}$, reflecting their sequence-level similarity. Homologous probability (H-Prob) reflects the probability that $p^{query}$ and $p^{target}$ is homologous.
A.3 More Results
Figure 6: (a) Distribution of pLDDT and SC-Perp. (b) Distribution of TM-score and Seq-Ident.
For the sequences generated without instructions by ProLLaMA, Figure 6 shows the distribution of pLDDT and SC-Perp, and Figure 6 shows the distribution of pLDDT and SC-Perp, where one spot denotes one sequence.

### v2

@@
-ProLLaMA: A Protein Large Language Model for Multi-Task Protein Language Processing
+ProLLaMA: A Protein Language Model for Multi-Task Protein Language Processing
 Abstract
-Large Language Models (LLMs), including GPT-x and LLaMA2, have achieved remarkable performance in multiple Natural Language Processing (NLP) tasks.
-Under the premise that protein sequences constitute the protein language, Protein Large Language Models (ProLLMs) trained on protein corpora excel at de novo protein sequence generation.
-However, as of now, unlike LLMs in NLP, no ProLLM is capable of multiple tasks in the Protein Language Processing (PLP) field.
-This prompts us to delineate the inherent limitations in current ProLLMs: (i) the lack of natural language capabilities, (ii) insufficient instruction understanding, and (iii) high training resource demands.
-To address these challenges, we introduce a training framework to transform any general LLM into a ProLLM capable of handling multiple PLP tasks.
-Specifically, our framework utilizes low-rank adaptation and employs a two-stage training approach, and it is distinguished by its universality, low overhead, and scalability.
-Through training under this framework, we propose the ProLLaMA model, the first known ProLLM to handle multiple PLP tasks simultaneously.
+Large Language Models (LLMs) have achieved remarkable performance in multiple Natural Language Processing (NLP) tasks.
+Under the premise that protein sequences constitute the protein language, Protein Language Models(PLMs) have advanced the field of protein engineering.
+However, as of now, unlike LLMs in NLP, PLMs cannot handle the protein understanding task and the protein generation task simultaneously in the Protein Language Processing (PLP) field.
+This prompts us to delineate the inherent limitations in current PLMs: (i) the lack of natural language capabilities, (ii) insufficient instruction understanding, and (iii) high training resource demands.
+To address these challenges, we introduce a training framework to transform any general LLM into a PLM capable of handling multiple PLP tasks.
+To improve training efficiency, we propose Protein Vocabulary Pruning (PVP) for general LLMs.
+We construct a multi-task instruction dataset containing 13 million samples with superfamily information, facilitating better modeling of protein sequence-function landscapes.
+Through these methods, we develop the ProLLaMA model, the first known PLM to handle multiple PLP tasks simultaneously.
 Experiments show that ProLLaMA achieves state-of-the-art results in the unconditional protein sequence generation task.
 In the controllable protein sequence generation task, ProLLaMA can design novel proteins with desired functionalities.
-In the protein property prediction task, ProLLaMA achieves nearly 100% accuracy across many categories.
-The latter two tasks are beyond the reach of other ProLLMs.
-Code is available at https://github.com/Lyu6PosHao/ProLLaMA.
-Machine Learning, ICML
+As for the protein understanding task, ProLLaMA achieves a 62% exact match rate in superfamily prediction.
+Codes, model weights, and datasets are available at https://github.com/PKU-YuanGroup/ProLLaMA and https://huggingface.co/GreatCaptainNemo.
 1 Introduction
-Large Language Models (LLMs), like GPT-x and LLaMA2 (Bubeck et al., 2023; Touvron et al., 2023), have achieved outstanding performance in handling a wide range of Natural Language Processing (NLP) tasks (Tamkin et al., 2021; Zhao et al., 2023; Kocoń et al., 2023; Qin et al., 2023; Huang et al., 2023; Zhong et al., 2023; Bang et al., 2023), including both Natural Language Generation (NLG) and Natural Language Understanding (NLU) tasks.
-This surge in LLMs has extended their applications beyond traditional contexts, including their adoption in the challenging field of protein design (Notin et al., 2022; Strokach & Kim, 2022; Ferruz & Höcker, 2022).
-Figure 1: Analogous to general LLMs managing multiple tasks in the NLP field, our ProLLaMA can handle multiple tasks in Protein Language Processing.
-In contrast, other ProLLMs are limited to the single task of unconditional protein generation.
-Taking protein sequences as the protein language, researchers train models with architectures similar to current LLMs on a vast protein corpus (Strokach & Kim, 2022; Ferruz et al., 2022; Nijkamp et al., 2023; Moffat et al., 2022; Madani et al., 2023; Qin et al., 2023).
-These Protein LLMs (ProLLMs) have paved the way for the rapid generation of structurally plausible protein sequences, holding immense potential for biomedical and biotechnological innovations (Notin et al., 2022; Nijkamp et al., 2023).
-However, this progress is met with challenges, particularly in extending their capabilities beyond sequence generation.
-Analogous to NLP, tasks related to protein language can be viewed as Protein Language Processing (PLP) (Bepler & Berger, 2021; Ofer et al., 2021).
-It is evident that current ProLLMs primarily focus on a single task in PLP, not covering multiple tasks like LLMs do in NLP (Ferruz et al., 2022; Madani et al., 2023).
-These limitations prompt the need for innovative solutions to unleash the full potential of ProLLMs.
-Developing a multi-tasking ProLLM would be highly beneficial for protein engineering and understanding the protein fitness landscape (Wright et al., 1932; Pan & Kortemme, 2021; Ren et al., 2022; Song & Li, 2023), but three main challenges must be considered:
-(i) Necessity of Natural Language: Protein language is not fully sufficient for PLP tasks, meaning it cannot fully represent all components of a task (the user instruction and the expected output) (Xu et al., 2023; Wang et al., 2023).
-The instruction and the output require a language beyond protein language (typically, natural language) for representation, which current ProLLMs lack.
-(ii) Instruction Following: To possess multi-tasking capabilities, models must execute tasks following user instructions (Zeng et al., 2023; Liu et al., 2023b; Zhou et al., 2023).
-However, current ProLLMs are unable to follow instructions.
-(iii) Training Resource Consumption: Substantial training resources are needed for models to learn natural language, protein language, and user instructions (Cui et al., 2023), which can sometimes be unaffordable.
-To address the challenges, we propose a two-stage training framework to achieve a ProLLM for multi-task PLP.
-In the first stage, we leverage a pre-trained general LLM like LLaMA2 to continually learn the protein language while maintaining the natural language knowledge of the model.
-In the second stage, the model is further trained on a multi-task PLP dataset through instruction tuning to equip the model with instruction following capabilities and multi-task capabilities of PLP.
-Notably, during both stages, we adopt Low-Rank Adaptation (LoRA) (Hu et al., 2021), which prevents catastrophic forgetting in the first stage, enhances scalability in the second stage and reduces the training cost.
-Using the training framework, we develop ProLLaMA, a model capable of multi-tasks for PLP, distinguishing it from all other ProLLMs.
-Through a series of experiments, we demonstrate the multitasking capabilities of our ProLLaMA.
-Specifically, in unconditional protein generation, ProLLaMA outperforms current ProLLMs and other sequence-generative models on common metrics such as pLDDT and TM-score.
-In controllable protein generation, based on a user-provided textual description of the desired protein functionalities, ProLLaMA generates novel proteins from scratch with functionalities as competent as natural proteins, such as SAM-MT and Trx.
-For protein property prediction, ProLLaMA achieves an average accuracy of 72% on the test dataset and nearly 100% accuracy in many subcategories.
+Large Language Models (LLMs), like GPT-x and LLaMA2 [1, 2], have achieved outstanding performance in handling a wide range of Natural Language Processing (NLP) tasks [3, 4, 5, 6, 7, 8, 9], including both Natural Language Generation (NLG) and Natural Language Understanding (NLU) tasks, in a generative manner.
+This surge in LLMs has extended their applications beyond traditional contexts, including their adoption in the challenging field of protein engineering [10, 11, 12, 13, 14].
+Taking protein sequences as the protein language, researchers train Protein Language Models (PLMs) on vast protein corpora [6, 11, 12].
+PLMs have the potential to significantly advance protein engineering, holding immense promise for biomedical and biotechnological innovations [10].
+However, this progress is challenged, particularly in extending their capabilities to multi-task Protein Language Processing (PLP).
+Analogous to NLP, tasks related to protein language can be viewed as PLP [15, 16].
+Consequently, PLP tasks can also be divided into two categories: Protein Language Generation (PLG), such as protein sequence generation, and Protein Language Understanding (PLU), such as protein property prediction.
+However, current PLMs tend to focus either exclusively on PLG [17, 18, 19] or PLU [20, 21, 22, 23, 24], rather than being proficient in both aspects simultaneously, as LLMs in NLP.
+These limitations prompt the need for innovative solutions to unleash the full potential of PLMs.
+Developing a multi-task PLM would be highly beneficial for protein engineering and protein fitness landscape modeling [25, 26, 27, 28], but three main challenges must be considered:
+(i) Necessity of Natural Language: Protein language is not fully sufficient for PLP tasks, meaning it cannot fully represent all components of a task (the task instruction, the input, and the expected output) [29, 30].
+It requires a language beyond protein language (typically, natural language) for representation, which current PLMs lack.
+(ii) Instruction Following: To possess multi-task capabilities, models must execute tasks following user instructions [31, 32, 33]. However, current PLMs are unable to follow instructions.
+(iii) Training Resource Consumption: Substantial training resources are needed for models to learn natural language, protein language, and user instructions [34], which can sometimes be unaffordable.
+To address the challenges, we construct an instruction dataset that contains approximately 13 million samples and encompasses both PLG and PLU tasks.
+We propose a two-stage training framework to achieve a PLM for multi-task PLP.
+In the first stage, we leverage a pre-trained general LLM like LLaMA2 to continually learn the protein language while maintaining the natural language knowledge.
+In the second stage, the model is further trained on the multi-task instruction dataset.
+Additionally, we propose Protein Vocabulary Pruning (PVP) for general LLMs to significantly improve training efficiency.
+Furthermore, during both stages, we adopt Low-Rank Adaptation (LoRA) [35], which prevents catastrophic forgetting in the first stage and reduces training costs in both stages.
+Figure 1: Left: LLMs can handle both generation and understanding tasks, whereas PLMs cannot.
+This highlights the disparity in capabilities between the two.
+Right: Our ProLLaMA can handle generation tasks (unconditional protein generation, controllable protein generation) and understanding tasks (protein superfamily prediction), surpassing current PLMs.
+Using these methods, we develop ProLLaMA, a model capable of multi-tasks for PLP, distinguishing it from all other PLMs.
+Through a series of experiments, we demonstrate the multi-task capabilities of our ProLLaMA.
+Specifically, as for PLG, in unconditional protein generation, ProLLaMA outperforms current PLMs on common metrics such as pLDDT and TM-score.
+In controllable protein generation, based on a user-provided textual description, ProLLaMA generates novel proteins from scratch with desired functionalities, such as the SAM-MT superfamily.
+As for PLU, in protein superfamily prediction, ProLLaMA achieved a 62% exact match rate on the test dataset and obtained an F1-score above 0.9 in many specific categories.
 In summary, the contributions of our research are as follows:
-$\bullet$ We propose a training framework with scalability and efficiency that enables any general LLM to be trained as a proficient model for multiple tasks in Protein Language Processing.
-$\bullet$ Our ProLLaMA is the first model to our knowledge capable of simultaneously handling multiple PLP tasks.
-$\bullet$ Experiments show that our ProLLaMA not only handles PLP tasks beyond the reach of existing ProLLMs but also achieves state-of-the-art results in the protein generation task where current ProLLMs are active.
+$\bullet$ We propose a training framework that enables any general LLM to be trained as a proficient model for multi-task PLP, including both PLG and PLU tasks.
+$\bullet$ We propose Protein Vocabulary Pruning and are the first to apply Low-Rank Adaptation in the training of large PLMs, which improves the efficiency of protein learning.
+$\bullet$ We construct an instruction dataset that contains 13 million samples and over 11,000 kinds of superfamily annotations, potentially facilitating a better modeling of sequence-function landscapes.
+$\bullet$ Experiments show that our ProLLaMA not only handles multiple PLP tasks but also achieves state-of-the-art results in protein generation tasks.
 2 Preliminaries
-2.1 Necessity of Natural Language
-As aforementioned, given the similarities between protein sequences and natural language, tasks related to protein sequences can be considered as Protein Language Processing (PLP), analogous to NLP.
+Figure 2: (A) Overview of the dataset construction.
+The protein language dataset contains 53 million samples, which is used for training in Stage 1.
+The instruction dataset contains 13 million instances with 11,268 unique superfamily annotations, which is used for training in Stage 2.
+(B) Overview of the training framework.
+Stage 1: The pre-trained LLaMA2 learns the protein language, resulting in ProLLaMA.
+Stage 2: ProLLaMA learns to perform multiple tasks by instruction tuning.
+(C) Overview of protein vocabulary pruning.
+The vocabulary, Embed layer and Head layer are pruned based on characteristics of protein datasets, aiming for higher training efficiency.
+Necessity of Natural Language.
+As aforementioned, given the similarities between protein sequences and natural language, tasks related to protein sequences can be considered as PLP, analogous to NLP.
 However, we observe a fundamental difference between protein language and natural language: natural language is complete for NLP tasks, whereas protein language is not complete for PLP tasks.
-Figure 2: The overall architecture of our ProLLaMA.
-We add low-rank adapters (LoRA) to certain weights.
-During the training process, we freeze these weights and other parameters, focusing solely on training LoRA.
-Specifically, natural language can represent all components of a task (i.e., the input instructions and the expected output) (Liu et al., 2023a).
-For example, in a sentiment analysis task, the instruction and output can be expressed as “Analyze the sentiment contained in this sentence: I am happy” and “The sentiment is positive” respectively, all in natural language.
-However, for PLP tasks, the instructions and outputs cannot be fully represented in protein language.
-For instance, in the protein property prediction task, the task instruction could be ”Predict the property of this protein: MAFCF…FEV”, with the expected output being ”The property is Trx superfamily.” It is evident that both the instructions and outputs of tasks require assistance from a language beyond protein language, in this case, natural language, for representation.
-Therefore, multi-task ProLLMs must possess a certain level of natural language ability, especially as more textual descriptions of proteins become available (Xu et al., 2023).
-2.2 Causal Language Modeling
-Causal Language Modeling (CLM) is the objective of training LLMs.
-Given a token sequence $\boldsymbol{x}=\{x_{0},x_{1},\ldots,x_{n-1}\}$ that is subject to training.
-CLM can be conceptualized as predicting the $i$-th token based on the preceding $i-1$ tokens (Bengio et al., 2000).
-Therefore, the optimization objective when training LLM using CLM is formulated as:
-$\small\mathcal{L}(\Theta)=\mathbb{E}_{\boldsymbol{x}\sim\mathcal{D}}\left[-\sum_{i}\log p(x_{i}|x_{0},x_{1},\ldots,x_{i-1};\Theta)\right]$ (1)
-where $\mathcal{L}$ denotes the loss function, $\Theta$ denotes the trainable parameters of the model, and $\mathcal{D}$ is the training dataset.
-2.3 Low Rank Adaptation
-Low-Rank Adaptation (LoRA) (Hu et al., 2021) is a parameter-efficient technique for fine-tuning LLMs.
+For example, in the sentiment analysis task, instructions, inputs, and outputs are straightforwardly expressed in natural language, such as “Analyze the sentiment: I am happy” yielding “The sentiment is positive.” However, in PLP tasks like protein property prediction, instructions like “Predict this protein’s property: MAFCF…FEV” cannot be fully conveyed in protein language alone, necessitating assistance from a language beyond protein language, in this case, natural language, for representation.
+Therefore, multi-task PLMs must possess a certain level of natural language ability, especially as more textual descriptions of proteins become available [29].
+Low-Rank Adaptation (LoRA) [35] is a parameter-efficient technique for fine-tuning LLMs.
 Due to the immense parameter size of LLMs, full-parameter fine-tuning could be impractical sometimes.
@@
 It achieves fine-tuning with a significantly smaller number of trainable parameters, yielding results comparable to full-parameter fine-tuning.
-[…]
-[…]
-[…]
-[…]
-Consequently, the quantity of parameters involved in training is reduced from $dh$ to $r(d+h)$.
-Given the low-rank hypothesis, where $r\ll d$ and $r\ll h$, this reduction in the number of trainable parameters is quite significant.
-Figure 3: The overall training framework of our ProLLaMA.
-(A-B): Training is divided into two stages.
-(A): By continual learning on UniRef50, LLaMA2 learns protein language, resulting in ProLLaMA.
-(B) By instruction tuning on the instruction dataset, ProLLaMA is capable of multi-tasking.
-(C) By instruction tuning again on your own dataset, ProLLaMA can be expanded to more tasks.
-During training, the original parameters of the model are frozen, with only the low-rank adapters trainable.
-After training, the newly acquired knowledge is stored in the low-rank adapters, namely A and B.
-Additionally, A and B can be integrated into the original model using Equation 2.
-Although this integration means the low-rank adapters are no longer plug-and-play, it ensures that the architecture of the post-training model remains consistent with that of the original model.
-Finally, LoRA prevents catastrophic forgetting of the original knowledge, as the newly learned knowledge has a lower rank than the original knowledge.
+LoRA prevents catastrophic forgetting of the original knowledge, as the newly learned knowledge has a lower rank than the original knowledge.
+The theoretical details of LoRA are provided in Appendix A.2.
 3 Methods
-The overview of the architecture of our ProLLaMA is shown in Figure 2, and the overview of our training framework is shown in Figure 3.
-In Section 3.1, we show how the LLaMA2 model learns protein language through continual learning, resulting in ProLLaMA.
-In Section 3.2, we show the integration of various tasks into ProLLaMA through instruction tuning.
-In Section 3.3, we show more tasks can be integrated into ProLLaMA.
-These three sections correspond to Figure 3(A-C).
-3.1 Continual Learning on Protein Language
-As mentioned in Section 1, current ProLLMs are lack of natural language abilities, which hinders multi-task capabilities.
-To solve this problem, we propose leveraging a pre-trained LLaMA2 to perform continual learning on protein language, as shown in Figure 3(A).
+Figure 3: The overview of the ProLLaMA model.
+We add low-rank adapters (LoRA) to certain weights and apply protein vocabulary pruning (PVP).
+We freeze original parameters, focusing solely on training LoRA (Embed and Head are also involved in the first training stage).
+In Section 3.1 and Figure 2(A), we show how to construct the protein language dataset and the instruction dataset.
+In Section 3.2 and Figure 2(C), we show how Protein Vocabulary Pruning (PVP) works.
+In Section 3.3, we show how the LLaMA2 model learns protein language on the protein language dataset through continued pre-training, resulting in ProLLaMA.
+In Section 3.4, we show the integration of various tasks into ProLLaMA through instruction tuning on the instruction dataset.
+The overview of the training framework is shown in Figure 2(B).
+The overview of the ProLLaMA model is shown in Figure 3.
+3.1 Dataset Construction
+The protein language dataset is utilized in the first training stage to enable LLaMA2 to grasp the language of proteins.
+Specifically, the dataset is sourced from UniRef50_2023_03 [36] on the UniProt website.
+We eliminate the descriptive parts of UniRef50, retaining only the pure protein sequences.
+Furthermore, We filter UniRef50 to ensure that the protein sequences consisted only of the 20 standard amino acids.
+We also retain sequences with a length of less than 512, aligning with ProGen [18].
+Given that the lengths of protein sequences follow a long-tail distribution, the sequences that are deleted constitute only a small portion of the total dataset.
+To preprocess the protein sequences, we employ a specific prefix “Seq=<” and suffix “>”.
+This standardized format aids LLaMA2 in distinguishing the new protein language from its existing natural language knowledge, thereby reducing confusion.
+The original Uniref50 contains 60,952,894 sequences, while after the above series of processing, our dataset comprises 52,807,283 protein sequences, with 90% for training and 10% reserved for testing.
+The instruction dataset is utilized in the second training stage to enable ProLLaMA to perform various tasks.
+We first obtain the protein2ipr database from InterPro [37], which includes all proteins from UniProtKB along with their corresponding InterPro annotation information.
+Subsequently, we iterate through each protein’s rep_id in UniRef50 to retrieve the corresponding annotation information from protein2ipr.
+This retrieval process is implemented using a distributed Redis database, and only proteins with lengths less than 256 participate in the retrieval to enhance efficiency.
+We utilize regular expressions to filter out superfamily annotation from the whole annotation.
+In the end, we obtain 6,350,106 data instances, each of which contains one protein sequence and its superfamily annotation.
+And the number of unique superfamily annotations is 11,268.
+Then, we process the obtained data into a multi-task instruction dataset following the Alpaca format [38], where each instance comprises three parts: instruction, input, and output.
+The instruction specifies the task type.
+We design two tasks: generating proteins based on superfamily and determining the superfamily of the given protein.
+For the former task, the input is the superfamily annotation, and the output is the expected protein.
+The latter task is the opposite.
+We selected these two tasks because they represent PLG and PLU, respectively.
+Additionally, modeling the relationships between sequences and superfamilies is crucial for uncovering protein functions and evolutionary insights [39].
+In the end, the instruction dataset comprises 12,700,212 (6,350,106 * 2) instances, with 90% for training and the rest reserved for testing.
+3.2 Protein Vocabulary Pruning
+It is known that LLMs are equipped with large vocabularies in tokenizers, which is beneficial to cover multi-lingual corpora and shorten the length of the input token sequence.
+However, a larger vocabulary also implies slower tokenization of raw text, along with increased parameters in the model’s Embed layer and Head layer, as discussed in detail in Appendix A.9.
+When considering only the protein language, such large vocabularies are clearly unnecessary.
+Therefore, we propose a method called Protein Vocabulary Pruning (PVP).
+PVP helps ProLLaMA achieve almost the same performance with higher training efficiency, whose proof is in Appendix A.3.
+Table 1: Effects of PVP. The vocabulary size, parameter numbers in Embed and Head layers, parameter numbers involved in training are listed.
+Training Stage Vocab Embed and Head Training
+I (w/o PVP) 32,000 262M 582M
+I (w/ PVP) 1,045 8.56M 328M
+II (w/o PVP) 32,000 262M 160M
+II (w/ PVP) 25,466 209M 160M
+PVP is rule-based.
+Specifically, we iterate through each token in the original vocabulary and retain it if it meets our predefined grammatical rules (e.g., consisting only of the 20 uppercase English letters representing standard amino acids).
+We save these tokens and their indices.
+Using the saved tokens, we construct a smaller vocabulary.
+We then use the indices to extract the corresponding vectors from Embed and Head layers to construct smaller ones.
+Thus, we can train LLMs using reduced vocabulary and parameters.
+PVP is employed before training.
+After training, the vocabulary can be recovered back to its original size by incorporating the entries from the original vocabulary.
+This process is also applicable to the Embed and Head layers, restoring them to their original dimensions.
+This ensures that the architecture of LLMs remains entirely consistent both before and after training, facilitating potential training on broader corpora in the future.
+Naturally, if subsequent training is not a consideration, omitting the recovery operation is permissible.
+We use PVP in both the first and second training stages of ProLLaMA, and its effects are shown in Table 1.
+It is evident that we achieve a significant compression rate in the first stage, primarily because the dataset in the first stage contains only protein sequences, which can be summarized by a simple rule.
+In contrast, the dataset in the second stage also includes English.
+More details about PVP is shown in Appendix A.3.
+3.3 Learning Protein Language
+As mentioned in Section 1, current PLMs lack natural language abilities, which hinders multi-task capabilities.
+To solve this problem, we propose leveraging LLaMA2 to perform continued pre-training on protein language.
 This approach is analogous to humans learning a foreign language, where the model learns protein language while retaining its original natural language abilities.
-Specifically, we construct a dataset based on UniRef50 (Suzek et al., 2015).
-We preprocess each protein sequence with specific prefixes and suffixes.
-This standardized format aids LLaMA2 in distinguishing the new protein language from its existing natural language knowledge, thereby reducing confusion.
-[…]
-We train the model with Equation 1 as the target on the aforementioned dataset, resulting in ProLLaMA.
-Benefiting from LoRA, we train only 10% of the parameters, in contrast to full-parameter training, which significantly reduces training costs.
-Additionally, as the remaining parameters are not involved in training, the inherent natural language abilities of the model are preserved.
-In summary, by enabling a pre-trained LLaMA2 to continually learn protein language and utilizing LoRA, we have developed ProLLaMA, a model that comprehends both protein sequence and natural language.
-Consequently, we have addressed the problem mentioned in the introduction: the lack of natural language abilities and excessive consumption of training resources.
-3.2 Performing Various Tasks
-As mentioned in Section 1, current ProLLMs are unable to perform multiple tasks based on user instructions.
-To solve this problem, we perform instruction tuning on the ProLLaMA obtained from the previous section, as shown in Figure 3(B).
-To be specific, we construct a multi-task dataset, where each item represents a task sample.
-A task consists of two parts: the instruction and the output.
-Training conducted on this dataset can be denoted as follows:
+We add Low-Rank Adapters (LoRA) into LLaMA2.
+To be specific, in each decoder block, we add LoRA to certain weights including $\textit{W}_{q}$, $\textit{W}_{k}$, $\textit{W}_{v}$, $\textit{W}_{o}$, $\textit{W}_{up}$, $\textit{W}_{gate}$ and $\textit{W}_{down}$.
+The original parameters are frozen, enabling only LoRA to be trained.
+Due to the significant differences between protein language and natural language, we choose a relatively high rank for LoRA, which helps the model learn protein sequences better and prevents under-fitting.
+We also include both the Embed and Head layers in training.
+This is based on the premise that a token may have different meanings in protein sequences and natural languages, requiring distinct embeddings for the same token.
+We trained the model using causal language modeling on the protein language dataset, resulting in ProLLaMA.
+Details on causal language modeling are provided in Appendix A.1.
+Benefiting from PVP and LoRA, as shown in Table 7, we train only about 5% of the parameters, in contrast to full-parameter training, which significantly reduces training costs.
+Additionally, as the remaining parameters are not involved in training, the inherent natural language abilities are preserved.
+In summary, we have developed ProLLaMA, a model that comprehends both protein language and natural language, with reduced training costs.
+Consequently, we have addressed two problems mentioned in Section 1: the lack of natural language abilities and excessive training costs.
+3.4 Performing Multiple Tasks
+As aforementioned, current PLMs are unable to perform multiple tasks based on user instructions. To solve this problem, we perform instruction tuning on ProLLaMA obtained from the previous section.
+We train ProLLaMA on the instruction dataset mentioned in Section 3.1:
 $\mathcal{L}(\Theta)=\mathbb{E}_{\boldsymbol{x,u}\sim\mathcal{D}}~{}[-\log~{}p(\boldsymbol{x}|\boldsymbol{u};\Theta)]$ (1)
 Here, $\Theta$ denotes the parameters to be optimized, $\mathcal{L}$ the loss function, and $\mathcal{D}$ the dataset.
-$\boldsymbol{u}$ denotes the instruction of a task.
-$\boldsymbol{x}=\{x_{0},x_{1},\ldots,x_{n-1}\}$ denotes the output of a task, where $x_{i}$ is the $i$-th token of the output.
-Since ProLLaMA employs causal language modeling, we need to combine Equation 4 with Equation 1:
-$\small\mathcal{L}(\Theta)=\mathbb{E}_{\boldsymbol{x}\sim\mathcal{D}}\left[-\sum_{i}\log p(x_{i}|\boldsymbol{u},x_{0},x_{1},\ldots,x_{i-1};\Theta)\right]$ (5)
+$\boldsymbol{u}$ denotes the instruction and the input of one instance.
+For brevity, when we refer to the instruction in the following text, it includes the input as well.
+$\boldsymbol{x}=\{x_{0},x_{1},\ldots,x_{n-1}\}$ denotes the output, where $x_{i}$ is the $i$-th token of the output.
+Since causal language modeling is employed, we need to combine Equation 1 with Equation 3:
+$\mathcal{L}(\Theta)=\mathbb{E}_{\boldsymbol{x}\sim\mathcal{D}}\left[-\sum_{i}\log p(x_{i}|\boldsymbol{u},x_{0},x_{1},\ldots,x_{i-1};\Theta)\right]$ (2)
 Equation 2 is the optimization objective for instruction tuning of ProLLaMA.
@@
 This is because the latter is the output part, where ensuring its quality of generation is crucial.
-The former, $\boldsymbol{u}$, as an instruction signal, only needs to be understood by the model and does not require generation.
+The former, $\boldsymbol{u}$, only needs to be understood by the model.
 In the instruction tuning stage, we exclusively train LoRA at a lower rank than specified in Section 3.3.
+In summary, through instruction tuning, we have made ProLLaMA capable of following instructions and performing multiple tasks.
+Consequently, we have addressed the problems mentioned in Section 1: the lack of instruction following and the lack of multi-task capabilities.
+In addition, the tasks that ProLLaMA can perform depend on the tasks included in the instruction dataset.
+Appendix A.7 shows that by training on an additional instruction dataset, ProLLaMA also performs well in predicting protein solubility.
 Table 2: Comparison of proteins generated by different models.
 Our ProLLaMA achieves the best performance on pLDDT, TM-score, and RMSD metrics, and is second-best in SC-Perp, demonstrating ProLLaMA excels in de novo protein design.
-Architecture Method pLDDTnormal-$\uparrow$ SC-Perpnormal-$\downarrow$ AFDB PDB
-TM-scorenormal-$\uparrow$ RMSDnormal-$\downarrow$ TM-scorenormal-$\uparrow$ RMSDnormal-$\downarrow$
-CNN CARP (Alamdari et al., 2023) 34.40±14.43 4.05±0.52 0.28 19.38 0.38 8.95
-LRAR (Alamdari et al., 2023) 49.13±15.50 3.59±0.54 0.40 14.47 0.43 9.47
-AutoEncoder ESM-1b (Rives et al., 2021) 59.57±15.36 3.47±0.68 0.34 20.88 0.44 8.59
-ESM-2 (Lin et al., 2023) 51.16±15.52 3.58±0.69 0.20 35.70 0.41 9.57
-Diffusion
-EvoDiff (Alamdari et al., 2023) 44.29±14.51 3.71±0.52 0.32 21.02 0.41 10.11
-LLM ProtGPT2 (Ferruz et al., 2022) 56.32±16.05 3.27±0.59 0.44 12.60 0.43 9.19
-ProGen2 (Nijkamp et al., 2023) 61.07±18.45 2.90±0.71 0.43 15.52 0.44 11.02
+*: We list ProGen2 and ProGen as the same item, referring to Appendix A.6 for explanation.
+AE: Auto-Encoder.
+AR: Auto-Regressive.
+[…]
+[…]
+CNN CARP [40] 34.40±14.43 4.05±0.52 0.28 19.38 0.38 8.95
+LRAR [40] 49.13±15.50 3.59±0.54 0.40 14.47 0.43 9.47
+PLM (AE) ESM-1b [22] 59.57±15.36 3.47±0.68 0.34 20.88 0.44 8.59
+ESM-2 [24] 51.16±15.52 3.58±0.69 0.20 35.70 0.41 9.57
+Diffusion EvoDiff [40] 44.29±14.51 3.71±0.52 0.32 21.02 0.41 10.11
+PLM (AR) ProtGPT2 [17] 56.32±16.05 3.27±0.59 0.44 12.60 0.43 9.19
+ProGen2/ProGen* [19, 18] 61.07±18.45 2.90±0.71 0.43 15.52 0.44 11.02
 ProLLaMA (ours) 66.49±12.61 3.10±0.65 0.49 9.50 0.48 7.63
-In summary, through instruction tuning, we have made ProLLaMA capable of following instructions and performing multiple tasks.
-In contrast, other ProLLMs remain limited to the single task of protein sequence generation.
-Consequently, we have addressed the problems mentioned in the introduction: the lack of instruction following and the lack of multi-task capabilities.
-3.3 Expanding to More Tasks
-In this section, we demonstrate that benefiting from the scalability of the training framework, ProLLaMA can be easily extended to a broader range of tasks as shown in Figure 3(C).
-To be specific, to adapt ProLLaMA to more tasks, researchers must customize the instruction dataset and then perform instruction tuning on ProLLaMA again with this dataset.
-ProLLaMA can be either the one obtained in the first stage or the one obtained in the second stage.
-Benefiting from LoRA, such instruction tuning requires only minimal training resources.
-An instance is shown in Figure 3(C).
-The new task in this instance is protein complex prediction, where the model must ascertain whether two input proteins can form a complex.
-In this task, the instruction $\boldsymbol{u}$ comprises a sentence describing the user intent and two protein sequences, while the output $\boldsymbol{x}$ should be a judgment of Yes or No.
-Table 2: Controllable generation on four different instructions.
-Given SAM-MT, TPHD, Trx, and CheY as instructions, ProLLaMA generates proteins with the desired functionalities.
-High values of TM-score and H-Prob indicate the generated proteins meet the instructions, in contrast to the uncontrollable generation of other models.
+Table 3: Controllable generation of ProLLaMA.
+SAM-MT, TPHD, Trx, and CheY are four supefamiles.
+High values of TM-score and H-Prob indicate the generated proteins meet the desired superfamily.
+“xxseen” denotes that xxresidues of one protein sequence belonging to the related superfamily are provided to ESM-1b for further generation.
+Even so, ProLLaMA performs best.
 Method SAM-MT TPHD Trx CheY
-TM-scorenormal-$\uparrow$ H-Probnormal-$\uparrow$ TM-scorenormal-$\uparrow$ H-Probnormal-$\uparrow$ TM-scorenormal-$\uparrow$ H-Probnormal-$\uparrow$ TM-scorenormal-$\uparrow$ H-Probnormal-$\uparrow$
-ESM-1b 0.58 0.37 0.55 0.48 0.61 0.37 0.63 0.27
-ESM-2 0.52 0.26 0.51 0.25 0.53 0.30 0.57 0.18
-EvoDiff 0.46 1.17 0.42 1.80 0.42 1.10 0.46 1.43
-ProtGPT2 0.45 3.86 0.43 4.62 0.44 2.53 0.45 4.86
-ProGen2 0.44 1.90 0.45 2.49 0.43 2.44 0.44 2.13
-ProLLaMA (ours) 0.71 98.13 0.82 100.00 0.93 99.96 0.81 100.00
-In summary, our framework and model are flexible and easily extensible. With collaborative efforts from other researchers, ProLLaMA has the potential to become more powerful.
+[…]
+ESM-1b 0.58 0.37% 0.55 0.48% 0.61 0.37% 0.63 0.27%
+ESM-2 0.52 0.26% 0.51 0.25% 0.53 0.30% 0.57 0.18%
+EvoDiff 0.46 1.17% 0.42 1.80% 0.42 1.10% 0.46 1.43%
+ProtGPT2 0.45 3.86% 0.43 4.62% 0.44 2.53% 0.45 4.86%
+ProGen2/ProGen* 0.44 1.90% 0.45 2.49% 0.43 2.44% 0.44 2.13%
+ESM-1b (25% seen) 0.43 0.64% 0.40 4.13% 0.42 1.01% 0.49 4.47%
+ESM-1b (50% seen) 0.59 61.08% 0.63 66.21% 0.64 62.75% 0.73 78.00%
+ESM-1b (75% seen) 0.67 88.51% 0.73 90.23% 0.75 93.92% 0.78 96.93%
+ProLLaMA (ours) 0.71 98.13% 0.82 100.00% 0.81 99.96% 0.93 100.00%
 4 Experiments
 We introduce the experiment setup in Section 4.1.
-And we evaluate the unconditional protein generation task in Section 4.2, the controllable protein generation task in Section 4.3, the protein property prediction task in Section 4.4, and the natural language ability in Section 4.5.
+And we evaluate the unconditional protein generation task in Section 4.2, the controllable protein generation task in Section 4.3, the protein property prediction task in Section 4.4.
 4.1 Experiment Setup
-Training Settings: For continual learning, the LoRA rank is set to 128, employing the AdamW optimizer alongside a cosine annealing scheduler with warm-up.
-The peak learning rate stands at 0.05, with a total of one training epoch.
-For instruction tuning, the LoRA rank is set to 64 with two training epochs, and all other settings remain consistent with the continual learning setup.
-Training Datasets: For continual learning, the dataset used consists of protein sequences from UniRef50 (Suzek et al., 2015).
-For instruction tuning, the instruction dataset is constructed using protein sequences from UniRef50 and protein property texts from InterPro (Paysan-Lafosse et al., 2023).
+Training Settings: For continued pre-training, the LoRA rank is set to 128, employing the AdamW optimizer alongside a cosine annealing scheduler with warm-up.
+The peak learning rate stands at 5e-5, with a total of one training epoch.
+It takes six days on eight A6000 GPUs using FlashAttention-2 [41].
+For instruction tuning, the LoRA rank is set to 64 with two training epochs, and all other settings remain consistent with the continued pre-training setup.
+It takes 5 days on eight A6000 GPUs.
+More training details can be found in Appendix A.4.
 Evaluation Settings: Unconditional protein generation involves generating protein sequences without specific instructions.
-Controllable protein generation involves generating desired protein sequences based on instructions that specify the required functionalities.
-Property prediction involves predicting protein properties based on instructions, which include the protein sequences to be predicted.
+Controllable protein generation involves generating desired protein sequences based on instructions that specify the required superfamily.
+Property prediction involves predicting protein superfamily based on instructions, which include the protein sequences to be predicted.
+All evaluations are conducted on one GPU with 24GB of VRAM, with model inference occupying approximately 13GB of VRAM.
 Evaluation Metrics: We use the following metrics to evaluate the generated protein sequences.
-The pLDDT (Jumper et al., 2021) is used to measure whether sequences are structurally plausible.
-Self-Consistency Perplexity (SC-Perp) (Alamdari et al., 2023) serves as an additional metric of plausible structures since pLDDT falls short in dealing with intrinsically disordered regions (IDRs) (Davey, 2019).
-TM-score (Zhang & Skolnick, 2004) reflects the structural similarity between generated sequences and known ones in AFDB (Varadi et al., 2022) and PDB (Berman et al., 2002).
+The pLDDT [42] is used to measure whether sequences are structurally plausible.
+Self-Consistency Perplexity (SC-Perp) [40] serves as an additional metric of plausible structures since pLDDT falls short in dealing with intrinsically disordered regions (IDRs) [43].
+TM-score [44] reflects the structural similarity between generated sequences and known ones in AFDB [45] and PDB [46].
 RMSD also reflects the structural similarity from the perspective of atomic distance.
@@
 Seq-Ident reflects the sequence similarity between generated sequences and known ones.
+More details are shown in Appendix A.5.
 4.2 Unconditional Protein Generation
-We compare our model with other state-of-the-art models in the protein sequence generation field.
-These models encompass a wide range of architectures and are trained exclusively on protein sequences.
+We compare our model with other models in protein sequence generation.
+These models cover a variety of types, with parameter numbers shown in Table 7.
 Table 2 shows the results.
@@
 This indicates that ProLLaMA, through its training on protein sequence data, can generate structurally plausible proteins.
-Additionally, it is important to note that ProLLaMA-generated proteins exhibit a mean and standard deviation for pLDDT and SC-Perp of 66.49±12.61 and 3.10±0.65, respectively.
-These values are comparable to those of natural proteins as reported by Alamdari et al.
-(2023), which are 68.25±17.85 and 3.09±0.63, respectively.
-Figure 4: Contrast experiment of ProLLaMA.
-(A-C): Compared to the baseline (ESM2), ProLLaMA maintains a high, and even higher, quality of generated proteins as their length increases.
-(D-F): For the four superfamilies of CheY, TPHD, Trx, and SAM-MT (SAM), the proteins generated by ProLLaMA are comparable to natural proteins.
-De novo design of long and structurally plausible protein sequences is highly challenging (Ferruz et al., 2022), yet our ProLLaMA excels in meeting this challenge.
-Figure 4(A) shows that as the sequence length increases, the pLDDT of proteins generated by ProLLaMA does not decrease but rather increases.
-In contrast, the pLDDT of proteins generated by ESM2 significantly decreases as the length increases.
-This indicates that ProLLaMA is able to capture long-range dependencies between amino acids, while ESM2 faces difficulties in the de novo design of long protein sequences.
-Figure 4(B) and Figure 4(C) also support this point.
-Figure 4(B) shows that as the length increases, SC-Perp of ProLLaMA slightly increases and then continues to decrease, while SC-Perp of ESM2 does the opposite.
-Figure 4(C) shows that as the length increases, the TM-score of ESM2 gradually decreases from above 0.6 to less than 0.2.
-In contrast, the TM-score of ProLLaMA remains above 0.5 for nearly all lengths, even exceeding 0.8 when the length is greater than 350.
-These demonstrate the robust sequence generation capability of ProLLaMA, especially in generating longer sequences.
+Notably, ProLLaMA-generated proteins exhibit a mean and standard deviation for pLDDT and SC-Perp of 66.49±12.61 and 3.10±0.65, respectively.
+These values are comparable to those of natural proteins as reported in [40], which are 68.25±17.85 and 3.09±0.63, respectively.
+It is noted that we list ProGen2 and ProGen as the same item in the table, referring to Appendix A.6 for explanation.
+Figure 4: Contrast experiments of ProLLaMA.
+(A-C): Compared to other models, ProLLaMA maintains a high quality of generated proteins as their length increases.
+(D-F): According to the three indicators normalized by percentage, proteins generated by ProLLaMA are roughly comparable to natural proteins in the four superfamilies of CheY, TPHD, Trx, and SAM-MT (SAM).
+De novo design of long and structurally plausible protein sequences is highly challenging [17], yet our ProLLaMA performs well.
+As shown in Figure 4(A)(B)(C), when the length is more than 300, ProLLaMA performs the best in all the three metrics.
+Although ProGen2’s performance is better in short sequences ($\leq$200), it decreases as the length increases.
+This indicates that ProLLaMA is able to capture long-range dependencies between amino acids while other models struggle.
 4.3 Controllable Protein Generation
-To instruct ProLLaMA in controllable protein generation, we utilize four superfamily descriptions as instructions respectively: the S-adenosyl-L-methionine-dependent methyltransferase superfamily (SAM-MT), the Tetratricopeptide-like helical domain superfamily (TPHD), the Thioredoxin-like superfamily (Trx), and the CheY-like superfamily (CheY).
+We utilize four superfamily descriptions as instructions respectively: the S-adenosyl-L-methionine-dependent methyltransferase superfamily (SAM-MT), the Tetratricopeptide-like helical domain superfamily (TPHD), the Thioredoxin-like superfamily (Trx), and the CheY-like superfamily (CheY).
 For each superfamily, ProLLaMA generates 100 protein sequences.
-Additionally, we use 100 natural proteins from each of the four superfamilies as benchmarks for comparison.
-We employ Foldseek to compare the generated proteins with the natural ones.
-Table 3: Protein property prediction. Our ProLLaMA achieves nearly 100% accuracy for many superfamilies. Here are 10 of them.
-OBFD UPF0145 NACD U3S CCHC Kazal SAM-MT TPHD Trx CheY
-Accuracy 100.00% 100.00% 100.00% 100.00% 95.24% 100.00% 93.67% 90.84% 94.17% 100.00%
+We randomly select 100 natural proteins from each of the four superfamilies as benchmarks for comparison.
+We employ Foldseek[47] to compare generated proteins with natural ones.
 The results shown in Table 3 demonstrate that ProLLaMA can generate desired protein sequences based on instructions that specify the required functionalities, confirming the capability for controllable generation.
@@
 In contrast, other models exhibit low TM-score and very low H-Prob due to their uncontrollable generation.
+We also conduct another experiment.
+We provide residues of natural proteins belonging to the superfamily to ESM-1b, allowing ESM-1b to complete these sequences, rather than generating them from scratch as done previously.
+As seen in Table 3, with more residues provided, the proteins generated by ESM-1b increasingly exhibit the characteristics of the corresponding superfamily.
+Even so, it still does not surpass ProLLaMA.
+Higher TM-score and H-prob indicate that, under instruction control, ProLLaMA’s de novo protein generation even outperforms the non-de novo generation by ESM-1b, which is provided with 75% of the residues.
+These results indicate that ProLLaMA can effectively capture structural and evolutionary relationships (as reflected by TM-score and H-prob) solely through learning from text and sequences.
 Additionally, using natural proteins as a benchmark, we assess pLDDT, SC-Perp, and TM-score of proteins generated by ProLLaMA.
@@
 Figure 4(F) visualizes the TM-score, with scores near the maximum indicating a high degree of structural similarity.
-These findings indicate that proteins generated by ProLLaMA are comparable to their natural counterparts in the same superfamily.
-Given that averages are considered, there are instances where generated proteins outperform natural proteins.
+These findings indicate that proteins generated by ProLLaMA are roughly comparable to their natural counterparts in the same superfamily.
+In Figure 5, we visualize four examples of proteins generated by ProLLaMA (colored in blue) alongside the most structurally similar natural proteins from PDB (colored in yellow).
+The significant overlap in 3D structures and the high TM-score confirm structural similarity.
+Low Seq-ident indicates sequence diversity.
+In summary, through controllable protein generation, ProLLaMA is capable of generating desired proteins with structures similar to natural proteins, yet with novel sequences.
+4.4 Property Prediction
+Table 4: Protein property prediction. The results of ten certain superfamilies in the test dataset.
+OBFD UPF0145 NACD U3S CCHC Kazal SAM-MT TPHD Trx CheY
+Precision 0.33 1.00 1.00 1.00 0.75 1.00 0.77 0.94 0.86 0.93
+Recall 1.00 1.00 1.00 1.00 0.95 1.00 0.94 0.91 0.94 1.00
+F1-score 0.50 1.00 1.00 1.00 0.86 0.98 0.84 0.93 0.90 0.96
 Figure 5: Protein visualization.
-Cases of controllable generation by ProLLaMA using SAM-MT, TPHD, Trx, and CheY as instructions.
+Four proteins of controllable generation by ProLLaMA using SAM-MT, TPHD, Trx, and CheY as instructions.
 Blue is generated proteins, and yellow is natural.
-They are similar in structure and function but different in sequence.
-In Figure 5, we present actual examples of proteins generated by ProLLaMA (depicted in blue) alongside the most structurally similar natural proteins from PDB (depicted in yellow).
-The four proteins are generated under the instructions of SAM-MT, TPHD, Trx, and CheY.
-We find that the four natural proteins also belong to SAM-MT, TPHD, Trx, and CheY, respectively, according to information from InterPro.
-This implies functional similarity between the generated proteins and natural ones, which further validates the effectiveness of controllable generation.
-The significant overlap in the 3D structures and the high TM-score confirm structural similarity.
-The lower Seq-Ident scores indicate sequence diversity.
-In summary, through controllable protein generation, ProLLaMA is capable of generating desired proteins with functions and structures similar to natural proteins, yet with novel sequences.
-4.4 Property Prediction
+They are similar in structure but different in sequence.
 We use the test dataset to evaluate whether ProLLaMA can predict the superfamily to which a given protein belongs.
 The test dataset consists of 10,000 samples.
-One protein may belong to two or more superfamilies simultaneously.
-Therefore, for the i-th protein in the test dataset, we denote its true superfamily text description as the set $F_{i}$ and the prediction by ProLLaMA as the set $F^{\prime}_{i}$.
-We calculate the prediction accuracy using:
-$\text{Acc}=\dfrac{\sum\nolimits_{i=1}^{N}\left|F_{i}\cap F^{\prime}_{i}\right|}{\sum\nolimits_{i=1}^{N}\left|F^{\prime}_{i}\right|}$ (6)
-where $\left|X\right|$ denotes the number of elements in set $X$, $\cap$ denotes the intersection of two sets, and $N$ denotes the number of samples in the test dataset.
-Although ProLLaMA actually performs a classification task here, it is more complex than typical ones.
+Although ProLLaMA performs a classification task here, it is more complex than typical ones.
 The key difference is that typical classification tasks require models to output a fixed label, often in one-hot encoding.
-In contrast, ProLLaMA outputs a full textual description of the result, significantly increasing the challenge.
-Even so, ProLLaMA still achieves an average accuracy of 72% on the test dataset.
-In addition, ProLLaMA achieves considerably high accuracy in many specific superfamily categories.
-Table 3 lists ten of them, showing that the accuracy exceeds 90% in all the ten superfamilies and reaches 100% in OBFD, UPF0145, NACD, U3S, Kazal, and CheY.
-4.5 Natural Language Ability
-Table 4: Comparison of the natural language ability. Vocab represents whether the vocabulary of the model supports natural language. Generation represents the sentence generation.
-Type Model Vocab Generation QA
-GeneralLLM LLaMA2 45% 44%
-ProLLM ProGen2 - -
-ProtGPT2 0% -
-ProLLaMA 26% 33%
-We evaluate the natural language ability of ProLLaMA and other models.
-As shown in Table 4, the vocabulary of ProGen2 only includes uppercase letters representing amino acids, failing to cover all English words and phrases.
-As for the sentence generation task on Wikipedia text, the accuracy of ProtGPT2 is 0%, indicating that it lacks natural language abilities.
-The accuracy of ProLLaMA is 26%, compared to 45% of LLaMA2.
-As for the QA task, the accuracy of ProLLaMA is 33%, compared to 44% of LLaMA2.
-These suggest that ProLLaMA has natural language abilities, albeit at a reduced level compared to LLaMA2.
+In contrast, ProLLaMA outputs the text.
+The advantage of the latter lies in its flexibility, such as the ability to easily handle situations where a sample belongs to multiple categories simultaneously.
+However, this increases task difficulty due to the much larger number of potential classification categories.
+Table 5: Protein property prediction. Results of all superfamilies in the test dataset.
+Metric Value
+Exact Match 0.62
+Jaccard Similarity 0.67
+Precision 0.63
+Recall 0.72
+F1-score 0.67
+Even so, as shown in Table 5, ProLLaMA generates superfamily descriptions that exactly match the real descriptions in 62% of the test dataset.
+In addition, Table 4 illustrates ProLLaMA’s performance on ten specific superfamilies.
+The recall value exceeds 0.9 in all the ten superfamilies and the F1-score exceeds 0.8 in nine superfamilies.
+The calculation formulas for these metrics can be found in Appendix A.5.
+Additional experiments on protein solubility prediction using ProLLaMA can be found in Appendix A.7.
 5 Related Work
-Protein Language Models
-ProLLMs aforementioned are a subset of Protein Language Models (PLMs).
-Recognizing the similarity between natural language sequences and protein sequences, many algorithms from NLP have been applied to protein sequence data (Yang et al., 2018; Alley et al., 2019; Rao et al., 2021; Elnaggar et al., 2021).
-This has led to the development of PLMs, which are broadly categorized into two types (Ferruz & Höcker, 2022; Zheng et al., 2023).
-One type adopts the decoder-only architecture and Causal Language Modeling (CLM) (Bengio et al., 2000; Vaswani et al., 2017), similar to general LLMs.
-Therefore, we refer to it as the Protein Large Language Model (ProLLM).
-ProLLMs primarily concentrate on de novo protein sequence generation (Moffat et al., 2022; Ferruz et al., 2022; Madani et al., 2023; Nijkamp et al., 2023), with a minority also focusing on fitness prediction (Notin et al., 2022).
-The other type adopts the encoder-only architecture and Masked Language Modeling (MLM) (Devlin et al., 2018; Meier et al., 2021; Rives et al., 2021; Brandes et al., 2022; Lin et al., 2023).
-They excel in protein representation learning, with the learned representations being applied to downstream tasks such as property prediction (Xu et al., 2023).
-However, they face challenges in de novo protein design.
-Our ProLLaMA is capable of multitasking, excelling in tasks that both of the above types specialize in, surpassing existing ProLLMs and even PLMs.
-Training LLMs
-There is a well-established procedure for training LLMs (Zhao et al., 2023).
-Initially, models undergo pre-training on large-scale, unlabeled corpora to grasp grammatical and semantic principles (Min et al., 2023).
-Then, models are subjected to instruction tuning using an instruction dataset, enabling models to understand instructions and perform various tasks (Zhang et al., 2023).
-However, current ProLLMs cannot undergo instruction tuning due to the lack of natural language abilities.
-Various parameter-efficient techniques have been proposed to accelerate training and conserve memory (He et al., 2021; Li & Liang, 2021; Hu et al., 2021; Liu et al., 2022; Hu et al., 2023), like Low-rank Adaptation (LoRA).
-We propose the use of LoRA in our training framework, which prevents catastrophic forgetting, enhances scalability, and reduces training costs.
-To our knowledge, this is the first application of parameter-efficient training techniques for ProLLMs.
+Protein Language Models.
+Recognizing the similarity between natural language sequences and protein sequences, many methods from NLP have been applied to protein sequence data [48, 49, 50, 51].
+This has led to the development of PLMs, which are broadly categorized into two types [12, 52]: Auto-Regressive (AR) PLMs and Auto-Encoder (AE) PLMs.
+AR PLMs adopt the decoder-only architecture and Causal Language Modeling (CLM) [53, 54].
+They primarily concentrate on PLG [55, 17, 18, 19], with a minority also focusing on fitness prediction [10].
+AE PLMs adopt the encoder-only architecture and Masked Language Modeling (MLM) [20, 21, 22, 23, 24].
+They excel in PLU, with the learned protein representations being applied to downstream predictive tasks [29].
+However, they face challenges in de novo protein generation.
+Our ProLLaMA is capable of multitasking, excelling in tasks that both of the above types specialize in, surpassing existing PLMs.
+This multitasking capability is achieved through instruction following, making it user-friendly.
+We have also noticed the recent emergence of scientific LLMs [13, 14, 56, 57, 58].
+In Appendix A.10, we discuss the differences between these models and ProLLaMA.
+Training LLMs.
+There is a general training framework for LLMs [4], where LLMs are first pre-trained on large-scale corpora [59] and then undergo instruction tuning to follow user instructions [60].
+However, we believe that our proposed two-stage training framework differs in motivation and insights, as discussed in Appendix A.11.
+Considering the vast number of parameters in LLMs, various parameter-efficient techniques have been proposed to accelerate training and conserve memory [61, 62, 35, 63, 64], including LoRA.
+For the same reason, some vocabulary pruning methods have been studied [65], which rely on statistics specific to a provided dataset.
+This results in them being more time-consuming and failing to leverage the prior patterns of the dataset.
+In contrast, our PVP, derived from the prior data patterns, achieves excellent compression rates.
 6 Conclusion
-Existing ProLLMs excel in the single task of protein generation but fall short in multiple tasks.
-In this work, we introduce an efficient training framework to transform any general LLM into a multi-task ProLLM.
-We also develop ProLLaMA, a versatile ProLLM for multiple tasks like controllable protein generation and protein property prediction.
+Existing PLMs excel in either protein generation tasks or protein understanding tasks.
+In this work, we introduce an efficient training framework to transform any general LLM into a multi-task PLM.
+We construct an instruction dataset containing both generation tasks and understanding tasks.
+We propose PVP to improve training efficiency.
+We develop ProLLaMA, a versatile PLM for multiple tasks like controllable protein generation and protein property prediction.
 Experiments indicate that ProLLaMA performs exceptionally well.
-We are confident that our framework and model will have a significant impact on the AI4Science community.
-Impact Statements
-Our work holds the potential to revolutionize the field of computational biology and biotechnology.
-With multi-task capabilities in the protein field, our ProLLaMA could significantly accelerate the pace of research and innovation in areas such as drug discovery, synthetic biology, and the development of novel biomaterials.
-Moreover, our work can enable a wider range of researchers and institutions, particularly those with limited resources, to participate in cutting-edge research, fostering a more inclusive AI4Science community.
-However, our work may raise concerns regarding model reliability and the potential for misuse.
-The uncertainty in model outputs could lead to misinformed decisions in critical research areas like drug discovery, necessitating cautious reliance and thorough validation by human experts to mitigate risks.
-Moreover, the powerful capabilities of ProLLaMA could be exploited for harmful purposes, posing bio-security challenges and ethical challenges.
-A collaborative effort among researchers, policymakers, and the broader community is crucial to harness the benefits of our work while addressing the potential risks and ethical considerations.
+We are confident that our work will have a significant impact on the AI4Science community.
 Appendix A Appendix
-A.1 Detailed Training Settings
-For continual learning, the block size is set to 2048, with a gradient accumulation step count of 8.
+A.1 Causal Language Modeling
+Causal Language Modeling (CLM) is a common objective of training LLMs.
+Given a token sequence $\boldsymbol{x}=\{x_{0},x_{1},\ldots,x_{n-1}\}$ that is subject to training.
+CLM can be conceptualized as predicting the $i$-th token based on the preceding $i-1$ tokens [53].
+Therefore, the optimization objective when training LLM using CLM is formulated as:
+$\mathcal{L}(\Theta)=\mathbb{E}_{\boldsymbol{x}\sim\mathcal{D}}\left[-\sum_{i}\log p(x_{i}|x_{0},x_{1},\ldots,x_{i-1};\Theta)\right]$ (3)
+where $\mathcal{L}$ denotes the loss function, $\Theta$ denotes the trainable parameters of the model, and $\mathcal{D}$ is the training dataset.
+A.2 Low-Rank Adaptation
+LoRA yields results comparable to full-parameter fine-tuning.
+Theoretically, fine-tuning can be conceptualized as a process of finding the change in parameters [66].
+Let the original parameters of the model be denoted as $\textit{W}_{0}$, and the parameters after fine-tuning as W.
+The objective of fine-tuning is to find $\Delta\textit{W}=\textit{W}-\textit{W}_{0}$.
+Hypothesizing $\Delta\textit{W}$ is of low rank [67], denoted as $r$, it can be decomposed into two low-rank matrices, $\Delta\textit{W}=\textit{AB}$.
+This leads to the following equation:
+$\textit{W}=\textit{W}_{0}+\textit{AB}$ (4)
+where $\textit{W},\textit{W}_{0}\in\mathbb{R}^{d\times h}$, $\textit{A}\in\mathbb{R}^{d\times r}$, and $\textit{B}\in\mathbb{R}^{r\times h}$.
+And the fine-tuning objective is transformed from finding $\Delta\textit{W}$ to finding $A$ and $B$:
+$\min_{\Delta\textit{W}}\mathcal{L}(\Delta\textit{W})\rightarrow\min_{A,B}\mathcal{L}(A,B)$ (5)
+Consequently, the quantity of parameters involved in training is reduced from $dh$ to $r(d+h)$.
+Given the low-rank hypothesis, where $r\ll d$ and $r\ll h$, this reduction in the number of trainable parameters is quite significant.
+Additionally, low-rank adapters can be integrated into the original model using Equation 4, which ensures that the architecture of the post-training model remains consistent with that of the original model.
+A.3 Protein Vocabulary Pruning
+Details.
+In the first training stage of ProLLaMA, the rule used by PVP is that the retained tokens consist only of uppercase letters representing the 20 standard amino acids, along with a few special tokens (e.g., <bos> token).
+In the second training stage, the rule used by PVP is that the retained tokens consist only of English letters, punctuation marks, and numbers, along with a few special tokens.
+After completion of the first stage training, the recovery operation is necessary to facilitate subsequent second stage training.
+The data reported in the experimental section are from models that do not use PVP.
+Proof of Concept.
+Casual language modeling is essentially a classification problem, with the classification space being identical to the input space, meaning the number of classes to classify is the size of the vocabulary, $V$.
+Suppose we know that the number of unique tokens present in the training set is $V^{\prime}$ (with $V^{\prime}$ being less than $V$).
+Thus, classes outside of $V^{\prime}$ are always negative classes during training for the model.
+In fact, PVP effectively incorporates this prior knowledge as an inductive bias into the model.
+This naturally does not harm the model’s training results on this training dataset.
+Therefore, for the first stage of training, the use of PVP does not affect the model’s performance.
+However, if we perform the PVP recovery operation and proceed to the subsequent second stage of training, the performance of the second stage training may be slightly inferior to that of the model trained without using PVP under the same settings.
+The following can compensate for this:
+During recovery, use the mean of the Embed layer obtained from the first stage model to fill the expanded Embed layer, and apply the same to the Head layer.
+In the second stage of training, add LoRA adapters to the Embed and Head layers and train them.
+A.4 Training Details
+For continued pre-training, the block size is set to 2048, with a gradient accumulation step count of 8.
 The warm-up ratio is 0.05, and the weight decay is set to 0.01.
@@
 DeepSpeed Zero Redundancy Optimizer stage 2 (ZeRO-2) is employed without using offload.
+FlashAttention-2 is also employed.
 For instruction tuning, the max sequence length is set to 256, with a gradient accumulation step count of 4.
@@
 The batch size per GPU is 144.
-Other settings are the same as continual learning.
-A.2 Detailed Evaluation Metrics
+Other settings are the same as continued pre-training.
+A.5 Detailed Evaluation Metrics
+A.5.1 Protein Generation
 We use OmegaFold to calculate the values of the predicted Local Distance Difference Test (pLDDT) of protein sequences.
@@
 Homologous probability (H-Prob) reflects the probability that $p^{query}$ and $p^{target}$ is homologous.
-A.3 More Results
-Figure 6: (a) Distribution of pLDDT and SC-Perp. (b) Distribution of TM-score and Seq-Ident.
-For the sequences generated without instructions by ProLLaMA, Figure 6 shows the distribution of pLDDT and SC-Perp, and Figure 6 shows the distribution of pLDDT and SC-Perp, where one spot denotes one sequence.
+A.5.2 Superfamily Prediction
+For the i-th protein in the test dataset, we denote its true superfamily text description as the set $F_{i}$ and the prediction by ProLLaMA as the set $F^{\prime}_{i}$.
+We calculate the prediction accuracy using:
+$\text{Recall}=\dfrac{\sum\nolimits_{i=1}^{N}\left|F_{i}\cap F^{\prime}_{i}\right|}{\sum\nolimits_{i=1}^{N}\left|F^{\prime}_{i}\right|}$
+$\text{Precision}=\dfrac{\sum\nolimits_{i=1}^{N}\left|F_{i}\cap F^{\prime}_{i}\right|}{\sum\nolimits_{i=1}^{N}\left|F_{i}\right|}$
+[…]
+[…]
+where $\left|X\right|$ denotes the number of elements in set $X$, $\cap$ denotes the intersection of two sets, and $N$ denotes the number of samples in the test dataset.
+The symbol $\delta(F_{i},F^{\prime}_{i})$ denotes the indicator function, which equals 1 when the sets $F_{i}$ and $F^{\prime}_{i}$ are exactly the same, and 0 otherwise.
+A.6 About ProGen2 and ProGen
+Following ProGen [18] and ProGen2 [19], it is known that ProGen enables controllable protein generation based on control tags, such as protein family, location, function, etc.
+And ProGen2 is limited to unconditional protein generation.
+Nevertheless, our attempts to find open-source model weights and executable code for ProGen were unsuccessful.
+In fact, the provided open-source repository link in the paper of ProGen is the same as that in the paper of ProGen2.
+However, the models provided by this repository cannot perform controllable protein generation.
+In addition, ProLLaMA can perform various tasks based on user instructions, including the protein understanding task, which ProGen cannot accomplish.
+A.7 Protein Solubility Prediction
+We transform the eSol dataset [68, 69] into an additional instruction dataset, which includes two tasks: generating proteins based on solubility and determining the solubility of proteins.
+We binarize the solubility, with “Solubility=<False>” indicating insoluble and “Solubility=<True>” indicating soluble.
+We train ProLLaMA (the one after the first training stage) on this additional instruction dataset.
+The LoRA rank is 64, the learning rate is 5e-5, and the number of training steps is 370.
+We compare our ProLLaMA with other methods in predicting protein solubility.
+The results are shown in Table 6.
+Our ProLLaMA outperforms models specifically designed for solubility prediction in terms of precision and F1-score.
+And its accuracy is almost the same as that of GraphSol.
+Notably, ProLLaMA utilizes only protein sequences, whereas other models incorporate additional features.
+Table 6: Protein Solubility Prediction. *:Values are sourced from GraphSol [70].
+Method Accuracy Precision Recall F1-score
+Protein-Sol* [71] 0.714 0.689 0.688 0.693
+DeepSol* [72] 0.763 0.771 0.738 0.695
+GraphSol* [70] 0.779 0.775 0.693 0.732
+ProLLaMA (ours) 0.775 0.788 0.685 0.733
+A.8 Comparison of Model Parameters
+Table 7 shows the comparison.
+As ProLLaMA takes a two-stage training procedure, 328M and 160M denote the parameters involved in Stage 1 and Stage 2, respectively.
+Due to the use of PVP and LoRA, ProLLaMA’s training parameters are significantly fewer than the inference parameters (which are also the total number of model parameters).
+In contrast, other models require full-parameter training, so their training parameters are equal to their inference parameters.
+ProGen2 has versions of different sizes; in this work, we compare against ProGen2-large.
+Table 7: Comparison of Model Parameters.
+CARP LRAR ESM-1b ESM-2 EvoDiff ProtGPT2 ProGen2 ProLLaMA
+Training parameters 640M 640M 650M 650M 640M 738M 2.7B (328+160)M
+Inference parameters 640M 640M 650M 650M 640M 738M 2.7B 7B
+A.9 Details about Embed layer and Head layer
+In fact, both layers can be viewed as "single-layer linear projection" layers.
+Embed layer projects tokens of the raw data to the embedding.
+The shape of it is 32000*4096, where 32000 is the size of the vocabulary, and 4096 is the feature size of the embedding.
+Head layer projects the embedding back to the language tokens space (probability distribution).
+By sampling from the probability distribution, the model can generate proteins and natural language.
+The shape of it is 4096*32000.
+Therefore, pruning the vocabulary can reduce the size of these two layers.
+A.10 Comparison with Scientific LLMs
+Recently, many scientific LLMs have emerged, addressing the deficiencies of conventional LLMs in specialized knowledge such as chemistry and biology, and also handling some protein-related tasks.
+However, we cannot classify them as PLMs because they have not been trained on large-scale protein corpora and have a limited understanding of the highly complex protein sequences.
+Additionally, their motivation is to supplement the shortcomings of LLMs in specific scientific domains.
+Therefore, they often use general LLMs, rather than PLMs, as baselines in experiment comparisons.
+Compared to PLMs, they perform poorly in protein-related tasks, especially in protein generation tasks.
+In contrast, the motivation behind ProLLaMA is to develop a more powerful PLM.
+As shown in Table 8, Mol-Instructions [13] performs worse than ProLLaMA and other PLMs in Table 2 on unconditional protein generation, which indicates that it struggles to generate structurally plausible proteins.
+As for controllable protein generation, Table 8 indicates that the control conditions in Mol-Instructions do not take effect.
+In summary, although scientific LLMs perform well in knowledge question answering tasks in fields such as medicine, biology, and chemistry, they do not master the protein language.
+Table 8: Comparison with scientific LLMs on unconditional protein generation.
+Type Method pLDDT $\uparrow$ TM-score $\uparrow$ RMSD $\downarrow$
+Scientific LLM Mol-Instructions 37.07±8.60 0.17 36.99
+PLM ProLLaMA 66.49±12.61 0.49 9.50
+Table 9: Comparison with scientific LLMs on controllable protein generation.
+Method SAM-MT TPHD Trx CheY
+[…]
+Mol-Instructions 0.39 6.55% 0.38 0.78% 0.39 1.29% 0.45 6.62%
+ProLLaMA 0.71 98.13% 0.82 100.00% 0.81 99.96% 0.93 100.00%
+A.11 Comparison with General Training Framework
+The general training framework of LLMs involves pre-training and instruction tuning.
+In fact, after pre-training, LLMs are already capable of performing multiple tasks due to the comprehensiveness of natural language as discussed in Section 2; instruction tuning is primarily for enhancing the quality of the output.
+In contrast, it is not feasible for LLMs to acquire multi-task PLP capability in the first stage due to the following factors: 1.
+Protein language is not complete for PLP tasks.
+2.
+Annotated protein sequences are significantly fewer than unannotated ones.
+3.
+The patterns within protein sequences are extremely complex, and LLMs do not have inherent knowledge of them.
+Therefore, we believe that the model should first learn the protein language in the initial stage, and in the second stage, learn to use both the protein language and natural language to perform PLP tasks.
+Additionally, our approach appears to have a certain biological plausibility.
+Specifically, our training framework can be viewed as modeling the patterns of protein sequences in the first stage and modeling the sequence-function landscape in the second stage.
+Compared to directly modeling the sequence-function landscape, our decoupled strategy significantly reduces the modeling difficulty.
+A.12 Impact Statements and Limitations
+Our work holds the potential to revolutionize the field of computational biology and biotechnology.
+With multi-task capabilities in the protein field, our ProLLaMA could significantly accelerate the pace of research and innovation in areas such as drug discovery, synthetic biology, and the development of novel biomaterials.
+Moreover, our work can enable a wider range of researchers and institutions, particularly those with limited resources, to participate in cutting-edge research, fostering a more inclusive AI4Science community.
+However, our work may raise concerns regarding model reliability and the potential for misuse.
+The uncertainty in model outputs could lead to misinformed decisions in critical research areas like drug discovery, necessitating cautious reliance and thorough validation by human experts to mitigate risks.
+Moreover, the powerful capabilities of ProLLaMA could be exploited for harmful purposes, posing bio-security challenges and ethical challenges.
+A collaborative effort among researchers, policymakers, and the broader community is crucial to harness the benefits of our work while addressing the potential risks and ethical considerations.
+A.13 Licenses and Guidelines
+As LLaMA2 is used in our work, all potential users of our ProLLaMA should respect the original license of Meta-LLaMA.
+Our datasets are based on UniProt and InterPro.
+UniProt applies CC BY 4.0 License.
+InterPro applies CC0 1.0 Public Domain Dedication.
+All potential users of our datasets should respect the licenses of UniProt and InterPro.
+We strongly advise all users to adhere to the highest ethical standards when using our model and datasets.
+This includes ensuring fairness, transparency, and responsibility in all aspects of their research and applications.
+Users should be vigilant in avoiding any forms of bias, discrimination, or unfair treatment that could arise from the use of our tools.
+It is crucial to maintain a commitment to accuracy and integrity, rigorously validating results and being transparent about methodologies.
+Any use of our model and datasets that could cause harm to individuals or society, whether through misinformation, invasion of privacy, or other unethical practices, is strictly prohibited.
+Users should contribute to a more responsible and beneficial use of AI technologies, promoting trust and positive impact in their respective fields.